### v1

# Random Polyhedral Cones: Distributional Results via Gale Duality

Abstract.
Let $U_{1},\ldots,U_{n}$ be independent random vectors uniformly distributed on the unit sphere $\mathbb{S}^{d-1}\subseteq\mathbb{R}^{d}$, where $n\geq d$, and consider the random polyhedral cone
$\mathcal{W}_{n,d}:=\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{n})=\{\lambda_{1}U_{1}+\ldots+\lambda_{n}U_{n}:\lambda_{1}\geq 0,\ldots,\lambda_{n}\geq 0\}.$
We establish several distributional results for $\mathcal{W}_{n,d}$ and the associated spherical polytope $\mathcal{W}_{n,d}\cap\mathbb{S}^{d-1}$. Our main contributions include:
  • (i)
    Let $\alpha_{d}$ denote the solid angle of $\mathcal{W}_{d,d}$ and write $m(d,k):=\mathbb{E}[\alpha_{d}^{k}]$ for its $k$-th moment. We prove the symmetry $m(d,k)=m(k,d)$. As an application, we compute $\mathop{\mathrm{Var}}\nolimits[\alpha_{d}]=2^{-d}(d+1)^{-1}-4^{-d}$ and derive a closed formula for the third moment.
  • (ii)
    For $n=d+1,d+2,d+3$ we determine the probability that $\mathcal{W}_{n,d}\cap\mathbb{S}^{d-1}$ is a spherical simplex, a spherical analogue of the classical Sylvester problem. In the case $n=d+2$ we also determine the distribution of the number of vertices of $\mathcal{W}_{d+2,d}\cap\mathbb{S}^{d-1}$.
  • (iii)
    Let $f_{\ell}(\mathcal{W}_{n,d})$ denote the number of $\ell$-dimensional faces of $\mathcal{W}_{n,d}$. We prove a distributional limit theorem for $f_{\ell}(\mathcal{W}_{n,d})$ in the regime $n=d+k$ and $\ell=d-q$, where $k,q\in\mathbb{N}$ are fixed and $d\to\infty$. The limit law is a weighted sum of independent chi squared variables, with weights given by explicit eigenvalues of a convolution operator on the sphere.
  • (iv)
    Let $I_{1},I_{2}\subseteq[n]$ be such that $I_{1}\cup I_{2}=[n]$ and $\max(\#I_{1},\#I_{2})\leq d-1$. We prove that the “face events” $\{\,\mathop{\mathrm{pos}}\nolimits(U_{i}:i\in I_{r})\text{ is a face of }\mathcal{W}_{n,d}\,\}$, $r=1,2$, are independent.
A unifying ingredient is an explicit coupling producing i.i.d. uniform vectors $U_{1},\ldots,U_{n}\in\mathbb{S}^{d-1}$ together with i.i.d. uniform vectors $V_{1},\ldots,V_{n}\in\mathbb{S}^{n-d-1}$ whose associated oriented matroids are Gale dual; faces of $\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{n})$ correspond to complementary subconfigurations of $V_{1},\ldots,V_{n}$ that positively span $\mathbb{R}^{n-d}$. This correspondence yields, in particular, a $U$-statistic representation of $f_{\ell}(\mathcal{W}_{n,d})$.
Key words and phrases:
Stochastic geometry, geometric probability, random polyhedral cones, random spherical polytopes, solid angles, Gaussian projection, Gale duality, spherical Sylvester problem, $U$-statistics, Funk–Hecke formula, spherical harmonics
2020 Mathematics Subject Classification:
Primary: 60D05, 52A22; Secondary: 52A55, 52B11, 52B35, 52B05, 33C55, 60F05
1. Introduction
1.1. Random polyhedral cones
Let $U_{1},\ldots,U_{n}$ be independent random vectors, each distributed uniformly on the unit sphere $\mathbb{S}^{d-1}$ in $\mathbb{R}^{d}$. The main object of interest in the present paper is the random polyhedral cone given by the positive hull of these vectors,
$\mathcal{W}_{n,d}:=\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{n}):=\{\lambda_{1}U_{1}+\cdots+\lambda_{n}U_{n}:\lambda_{1}\geq 0,\ldots,\lambda_{n}\geq 0\}\subseteq\mathbb{R}^{d}.$
The polar (convex dual) cone of $\mathcal{W}_{n,d}$ is the intersection of the half-spaces with outer unit normal vectors $U_{1},\ldots,U_{n}$, namely
$\mathcal{W}_{n,d}^{\circ}:=\{y\in\mathbb{R}^{d}:\langle x,y\rangle\leq 0\text{ for all }x\in\mathcal{W}_{n,d}\}=\bigcap_{i=1}^{n}\{y\in\mathbb{R}^{d}:\langle y,U_{i}\rangle\leq 0\}.$
Thus, $\mathcal{W}_{n,d}^{\circ}$ is the feasible set of a system of $n$ random linear inequalities, making it a natural and fundamental object. The random cones $\mathcal{W}_{n,d}$, $\mathcal{W}_{n,d}^{\circ}$, and related models have been studied by Cover and Efron [8], Donoho and Tanner [11], and Hug and Schneider [20]; see also [15, 21, 22, 25, 41] and Chapters 5–6 of the book by Schneider [42]. One important result is Wendel’s formula [47] (see also [43, Theorem 8.2.1]), which states that
(1.1) $p(n,d):=\mathbb{P}[\mathcal{W}_{n,d}=\mathbb{R}^{d}]=\frac{1}{2^{n-1}}\sum_{\ell=d}^{n-1}\binom{n-1}{\ell}$.
Beyond this, explicit formulas are known for expectations of several geometric functionals of $\mathcal{W}_{n,d}$ and $\mathcal{W}_{n,d}^{\circ}$ (as well as of their conditioned versions). As one example, for $\ell\in\{0,\ldots,d-1\}$, let $f_{\ell}(\mathcal{W}_{n,d})$ denote the number of $\ell$-dimensional faces of $\mathcal{W}_{n,d}$. Then Donoho and Tanner [11, Theorem 1.6] show that
$\mathbb{E}f_{\ell}(\mathcal{W}_{n,d})=\binom{n}{\ell}\,p(n-\ell,n-d)=\frac{1}{2^{\,n-\ell-1}}\binom{n}{\ell}\sum_{p=n-d}^{n-\ell-1}\binom{n-\ell-1}{p}$.
Other geometric functionals for which explicit expectation formulas are available include the conic intrinsic volumes (with the solid angle $\alpha(\mathcal{W}_{n,d})$ as a special case), as well as sums of conic intrinsic volumes over all faces of a given dimension and sums over the tangent cones at these faces. There also exist sporadic results on second moments of certain geometric functionals [20, Theorem 8.1], but almost nothing seems to be known about the full distribution of $f_{\ell}(\mathcal{W}_{n,d})$, $\alpha(\mathcal{W}_{n,d})$, and related quantities.
1.2. Main results
In the present paper we establish several distributional results for $\mathcal{W}_{n,d}$ that go beyond expectations and second moments.
Moments of the solid angle
In Section 4 we study the solid angle of the cone $\mathcal{W}_{d,d}$ (i.e., $n=d$). For a full-dimensional polyhedral cone $C\subseteq\mathbb{R}^{d}$, its solid angle is defined by $\alpha(C):=\mathbb{P}[U\in C]$, where $U$ is uniformly distributed on the unit sphere in $\mathbb{R}^{d}$. Writing $m(d,k):=\mathbb{E}[\alpha(\mathcal{W}_{d,d})^{k}]$ for the $k$-th moment, we prove the symmetry relation $m(d,k)=m(k,d)$. As an application, we compute
$\mathop{\mathrm{Var}}\nolimits(\alpha(\mathcal{W}_{d,d}))=2^{-d}(d+1)^{-1}-4^{-d}.$
We also derive a closed-form expression for the third moment and propose a conjecture for the asymptotic behavior of $m(d,k)$ as $d\to\infty$ (with fixed $k$).
Spherical Sylvester problem
In Section 5 we address a spherical analogue of Sylvester-type questions. For $n=d+1,d+2,d+3$ we compute the probability that $\mathcal{W}_{n,d}\cap\mathbb{S}^{d-1}$ is a spherical simplex. Moreover, for $n=d+2$ we determine the distribution of the number of vertices of $\mathcal{W}_{d+2,d}\cap\mathbb{S}^{d-1}$.
High-dimensional limit theorems for face counts
In Section 6 we prove a distributional limit theorem for face counts in a high-dimensional regime. Recall that $f_{\ell}(\mathcal{W}_{n,d})$ denotes the number of $\ell$-dimensional faces of $\mathcal{W}_{n,d}$. We consider $n=d+k$ and $\ell=d-q$, where $k\in\mathbb{N}$ and $q\in\mathbb{N}$ are fixed, and let $d\to\infty$. We show that the centered and normalized quantity $f_{d-q}(\mathcal{W}_{d+k,d})$ converges in distribution to a non-Gaussian limit; more precisely,
$d\cdot\left(\frac{f_{d-q}(\mathcal{W}_{d+k,d})}{\binom{d+k}{d-q}}-p(k+q,k)\right)\;{\overset{w}{\underset{d\to\infty}{\longrightarrow}}}\;\frac{\binom{k+q}{2}\binom{k+q-2}{k-1}}{2^{k+q-1}}\,(1-Q_{k}),$
where $Q_{k}>0$ is an infinitely divisible random variable with $\mathbb{E}Q_{k}=1$ that admits the representation
$Q_{k}\stackrel{{\scriptstyle d}}{{=}}\sum_{r=1,3,5,\ldots}\left(\frac{\Gamma\!\left(\frac{k}{2}\right)\Gamma\!\left(\frac{r}{2}\right)}{\pi\,\Gamma\!\left(\frac{r+k}{2}\right)}\right)^{2}\operatorname{Gamma}\left(\frac{d_{r,k}}{2},\frac{1}{2}\right),\qquad k\geq 2,$
and $Q_{1}\stackrel{{\scriptstyle d}}{{=}}\operatorname{Gamma}(\tfrac{1}{2},\tfrac{1}{2})$. Here $\operatorname{Gamma}(\alpha,\lambda)$ denotes a Gamma random variable with shape parameter $\alpha>0$ and rate parameter $\lambda>0$, and all such variables are independent. Moreover,
$d_{r,k}:=\binom{r+k-1}{r}-\binom{r+k-3}{r-2}$
is the dimension of the space of degree-$r$ spherical harmonics on $\mathbb{S}^{k-1}\subseteq\mathbb{R}^{k}$ for $k\geq 2$. A more detailed statement is given in Theorem 6.1.
Independence of face events
For a subset $I\subseteq\{1,\ldots,n\}$, define the cone $F_{I}:=\mathop{\mathrm{pos}}\nolimits(U_{i}:i\in I)$, which is a candidate face of $\mathcal{W}_{n,d}$, and the event
$A_{I}:=\{F_{I}\text{ is a face of }\mathcal{W}_{n,d}\}.$
Somewhat surprisingly, many of the “face events” $A_{I}$ are independent. In Section 3.1 we provide a sufficient condition guaranteeing such independence.
1.3. Common technique: Gale duality
Although the results above may appear unrelated at first sight, their proofs share a common ingredient: passing to the (linear) Gale transform. The necessary background on Gale transforms is collected in Section 2.1.
The present work was inspired by Frick et al. [14]. Their key observation is that the affine Gale transform of $n$ i.i.d. standard Gaussian points in $\mathbb{R}^{d}$ may be realized as $n$ independent standard Gaussian points in $\mathbb{R}^{\,n-d-1}$, translated so that their barycenter is at the origin. Using this, they established an equivalence between the following two problems:
  • (a)
    determining the distribution of the Radon type of $d+2$ independent Gaussian points in $\mathbb{R}^{d}$, and
  • (b)
    determining the location of the sample mean among the order statistics of a one-dimensional i.i.d. Gaussian sample of size $d+2$.
For further results on these two problems, see [4, 7, 23, 28, 48].
In contrast, the Gale duality input used in the present paper is linear rather than affine. The corresponding results will be stated in Proposition 2.5 and Theorem 2.8.
1.4. Notation
Throughout the paper we use the following notation. For a finite set $A$, let #A\#A denote its cardinality. For $n\in\mathbb{N}$, write $n]\coloneqq\{1,\ldots,n\}$.
Let $\mathbb{R}^{d}$ be $d$-dimensional Euclidean space with the standard inner product $\langle\cdot,\cdot\rangle$ and norm $\|\cdot\|$. Let $\mathbb{S}^{d-1}:=\{x\in\mathbb{R}^{d}:\|x\|=1\}$ be the unit sphere, and let $e_{1},\ldots,e_{d}$ denote the standard orthonormal basis of $\mathbb{R}^{d}$.
For $X\subseteq\mathbb{R}^{d}$, we write $\mathop{\mathrm{pos}}\nolimits(X)$ for the positive hull of $X$ and $\mathop{\mathrm{lin}}\nolimits(X)$ for the linear span of $X$, that is,
$\mathop{\mathrm{pos}}\nolimits(X)$ $:=\left\{\sum_{i=1}^{m}\lambda_{i}x_{i}\,:\,m\in\mathbb{N}_{0},\ x_{1},\ldots,x_{m}\in X,\ \lambda_{1},\ldots,\lambda_{m}\geq 0\right\},$
$\mathop{\mathrm{lin}}\nolimits(X)$ $:=\left\{\sum_{i=1}^{m}\lambda_{i}x_{i}\,:\,m\in\mathbb{N}_{0},\ x_{1},\ldots,x_{m}\in X,\ \lambda_{1},\ldots,\lambda_{m}\in\mathbb{R}\right\}.$
If $X=\{x_{1},\ldots,x_{n}\}$ is finite, we also write $\mathop{\mathrm{pos}}\nolimits(x_{1},\ldots,x_{n})$ and $\mathop{\mathrm{lin}}\nolimits(x_{1},\ldots,x_{n})$ instead of $\mathop{\mathrm{pos}}\nolimits(X)$ and $\mathop{\mathrm{lin}}\nolimits(X)$. By convention, $\mathop{\mathrm{pos}}\nolimits(\varnothing)=\mathop{\mathrm{lin}}\nolimits(\varnothing)=\{0\}$.
A polyhedral cone is the positive hull of finitely many vectors in $\mathbb{R}^{d}$. Equivalently, it is an intersection of finitely many half-spaces of the form $\{x\in\mathbb{R}^{d}:\langle x,u\rangle\leq 0\}$, where $u\in\mathbb{R}^{d}\setminus\{0\}$.
For sequences $(a_{n})_{n\in\mathbb{N}}$ and $(b_{n})_{n\in\mathbb{N}}$, we write $a_{n}\sim b_{n}$ if $a_{n}/b_{n}\to 1$ as $n\to\infty$. For random variables (or random elements) $X$ and $Y$, we write $X\stackrel{{\scriptstyle d}}{{=}}Y$ to denote equality in distribution. Weak convergence of random variables or elements is denoted by $\overset{w}{\longrightarrow}$.
2. Linear Gale duality for random vector configurations
2.1. Linear Gale duality
Gale duality (or the Gale transform) is a widely used tool in convex and discrete geometry. For detailed accounts of Gale duality and its applications, we refer to the books of Grünbaum [17, Section 5.4], Ziegler [50, Chapter 6],  Matoušek [33, Section 5.6], McMullen and Shephard [35, Chapter 3], Björner et al. [6, Section 3.4], and to the papers of McMullen [34], Shephard [46], and Grünbaum and Shephard [18]. There are two versions of Gale duality: affine and linear. The version most commonly treated in the literature is the affine one; in the present work, however, we require linear Gale duality studied in [34, 46], which we now define.
Definition 2.1 (Linear Gale duality).
Let $d^{\prime},d^{\prime\prime}\in\mathbb{N}$ and put $n:=d^{\prime}+d^{\prime\prime}$. Two vector configurations $a_{1},\dots,a_{n}\in\mathbb{R}^{d^{\prime}}$ and $b_{1},\dots,b_{n}\in\mathbb{R}^{d^{\prime\prime}}$ are said to be in linear Gale duality if the following conditions hold:
  1. (i)
    the linear span of $a_{1},\dots,a_{n}$ is $\mathbb{R}^{d^{\prime}}$ and the linear span of $b_{1},\dots,b_{n}$ is $\mathbb{R}^{d^{\prime\prime}}$;
  2. (ii)
    if $A\in\mathbb{R}^{d^{\prime}\times n}$ is the matrix with comumns $a_{1},\ldots,a_{n}$ and $B\in\mathbb{R}^{d^{\prime\prime}\times n}$ is the matrix with columns $b_{1},\ldots,b_{n}$, then every row of $A$ is orthogonal to every row of $B$. In matrix notation,
    $AB^{\top}\;=\;0_{d^{\prime}\times d^{\prime\prime}}.$
Condition (i) states that $\mathop{\mathrm{rank}}\nolimits A=d^{\prime}$ and $\mathop{\mathrm{rank}}\nolimits B=d^{\prime\prime}$. The row spaces of $A$ and $B$, denoted by $\operatorname{Row}(A)$ and $\operatorname{Row}(B)$, are defined as the linear subspaces in $\mathbb{R}^{n}$ spanned by the rows of $A$ and $B$, respectively. Condition (ii) states that $\operatorname{Row}(A)$ and $\operatorname{Row}(B)$ are orthogonal to each other. In fact, by the full rank assumption, they even form complementary orthogonal subspaces of $\mathbb{R}^{n}$. Thus conditions (i) and (ii) can be stated as
(2.1) $\dim\operatorname{Row}(A)=d^{\prime},\qquad\dim\operatorname{Row}(B)=d^{\prime\prime},\qquad\operatorname{Row}(A)=\operatorname{Row}(B)^{\perp}.$
There is a pronounced correspondence between the properties of linearly Gale-dual configurations $a_{1},\dots,a_{n}$ and $b_{1},\dots,b_{n}$. The following Lemma 2.2 records a concrete instance of this correspondence: faces of the cone generated by the $a_{i}$ correspond to complements of positive spanning subsets of the $b_{j}$, and conversely. While this fact is standard and can be found in some or other form in the references cited above, we include a complete elementary proof in Appendix A, Lemma A.4.
Lemma 2.2 (Faces vs. positive spanning subsets under general linear position).
Let $d^{\prime},d^{\prime\prime}\in\mathbb{N}$ and put $n:=d^{\prime}+d^{\prime\prime}$. Let $a_{1},\dots,a_{n}\in\mathbb{R}^{d^{\prime}}$ and $b_{1},\dots,b_{n}\in\mathbb{R}^{d^{\prime\prime}}$ be in linear Gale duality, and assume that both configurations are in general linear position. Let $I\subseteq[n]$ satisfy $I\leq d^{\prime}-1$. Then the following are equivalent:
  1. (i)
    $\mathop{\mathrm{pos}}\nolimits(a_{i}:i\in I)$ is a face of the polyhedral cone $\mathop{\mathrm{pos}}\nolimits(a_{1},\dots,a_{n})$.
  2. (ii)
    $\mathop{\mathrm{pos}}\nolimits(b_{j}:j\in I^{c})=\mathbb{R}^{d^{\prime\prime}}$. Here $\II^{c}:=[n]\backslash I$ is the complement of $I$.
Remark 2.3.
Pairwise distinct vectors $v_{1},\dots,v_{m}\in\mathbb{R}^{d}$ are said to be in general linear position if every subset of $\{v_{1},\dots,v_{m}\}$ of size at most $d$ is linearly independent. If $m\geq d$, this is equivalent to requiring that every $d$-element subset of $\{v_{1},\dots,v_{m}\}$ is linearly independent.
Example 2.4.
Let $I=\varnothing$. By convention, $\mathop{\mathrm{pos}}\nolimits(\varnothing)=\{0\}$, so condition (i) says that $\{0\}$ is a face of $C_{A}:=\mathop{\mathrm{pos}}\nolimits(a_{1},\ldots,a_{n})$, i.e. that the cone $C_{A}$ is pointed. If the vectors $a_{1},\ldots,a_{n}$ are in general linear position and linearly span $\mathbb{R}^{d^{\prime}}$, then one checks that $C_{A}$ is either pointed or equal to $\mathbb{R}^{d^{\prime}}$. Hence Lemma 2.2 with $I=\varnothing$ takes the form
$\mathop{\mathrm{pos}}\nolimits(a_{1},\ldots,a_{n})\neq\mathbb{R}^{d^{\prime}}\quad\Longleftrightarrow\quad\mathop{\mathrm{pos}}\nolimits(b_{1},\ldots,b_{n})=\mathbb{R}^{d^{\prime\prime}}.$
2.2. Gale coupling between two Gaussian projections
Let $\mathbf{G}$ be a Gaussian $d_{1}\times d_{2}$ matrix, i.e. a random matrix whose entries are i.i.d. standard Gaussian random variables. We view $\mathbf{G}$ as a random linear operator $\mathbf{G}:\mathbb{R}^{d_{2}}\to\mathbb{R}^{d_{1}},$ and refer to it as a Gaussian linear operator. Its transpose $\mathbf{G}^{\top}:\mathbb{R}^{d_{1}}\to\mathbb{R}^{d_{2}}$ is again a Gaussian linear operator. A key property is rotational invariance: for any orthogonal transformations $O_{1}\in O(d_{1})$ and $O_{2}\in O(d_{2})$, the random matrix $O_{1}\,\mathbf{G}\,O_{2}$ has the same distribution as $\mathbf{G}$. In particular, the notion of a Gaussian linear operator does not depend on the choice of orthonormal bases in $\mathbb{R}^{d_{1}}$ and $\mathbb{R}^{d_{2}}$. If $d_{2}\geq d_{1}$, we also refer to $\mathbf{G}$ as a Gaussian projection; note, however, that $\mathbf{G}$ is not a projection in the literal sense.
Fix $d^{\prime},d^{\prime\prime}\in\mathbb{N}$ and set $n:=d^{\prime}+d^{\prime\prime}$. Let $\mathbf{G}:\mathbb{R}^{n}\to\mathbb{R}^{d^{\prime}}$ and $\mathbf{H}:\mathbb{R}^{n}\to\mathbb{R}^{d^{\prime\prime}}$ be Gaussian projections. With probability one, both $\operatorname*{Ker}(\mathbf{G})$ and $\operatorname*{Ker}(\mathbf{H})^{\perp}$ are $d^{\prime\prime}$-dimensional linear subspaces of $\mathbb{R}^{n}$ (since $\mathbf{G}$ has rank $d^{\prime}$ and $\mathbf{H}$ has rank $d^{\prime\prime}$ almost surely). Let $G(n,d^{\prime\prime})$ denote the Grassmannian of $d^{\prime\prime}$-dimensional linear subspaces of $\mathbb{R}^{n}$, endowed with its natural Borel $\sigma$-algebra. We view $\operatorname*{Ker}(\mathbf{G})$ and $\operatorname*{Ker}(\mathbf{H})^{\perp}$ as random elements of $G(n,d^{\prime\prime})$. By rotational invariance of Gaussian matrices, these random subspaces are $O(n)$-invariant: for every orthogonal transformation $O\in O(n)$,
$O\,\operatorname*{Ker}(\mathbf{G})\ \stackrel{{\scriptstyle d}}{{=}}\ \operatorname*{Ker}(\mathbf{G})\qquad\text{and}\qquad O\,\operatorname*{Ker}(\mathbf{H})^{\perp}\ \stackrel{{\scriptstyle d}}{{=}}\ \operatorname*{Ker}(\mathbf{H})^{\perp}.$
Since there is a unique $O(n)$-invariant probability measure on $G(n,d^{\prime\prime})$ (the Haar, or uniform, measure), it follows that both $\operatorname*{Ker}(\mathbf{G})$ and $\operatorname*{Ker}(\mathbf{H})^{\perp}$ are distributed according to this measure. So
(2.2) $\operatorname*{Ker}(\mathbf{G})\ \stackrel{{\scriptstyle d}}{{=}}\ \operatorname*{Ker}(\mathbf{H})^{\perp},\qquad\text{and this common distribution is uniform on }G(n,d^{\prime\prime}).$
The next proposition shows that one can couple two Gaussian projections $\mathbf{G}$ and $\mathbf{H}$ on a common probability space in such a way that the distributional identity (2.2) is realized by an almost sure equality of subspaces. ($\mathbf{G}$ and $\mathbf{H}$ may become dependent in this coupling.)
Proposition 2.5 (Coupling of Gaussian projections).
Let $d^{\prime},d^{\prime\prime}\in\mathbb{N}$ and set $n:=d^{\prime}+d^{\prime\prime}$. There exists a probability space on which one can define two Gaussian projections $\mathbf{G}:\mathbb{R}^{n}\to\mathbb{R}^{d^{\prime}}$ and $\mathbf{H}:\mathbb{R}^{n}\to\mathbb{R}^{d^{\prime\prime}}$ such that
(2.3) $\operatorname*{Ker}\mathbf{G}\;=\;(\operatorname*{Ker}\mathbf{H})^{\bot}\quad\text{a.s.}$
Moreover, if $e_{1},\ldots,e_{n}$ denotes the standard orthonormal basis of $\mathbb{R}^{n}$, then for almost every realization of $(\mathbf{G},\mathbf{H})$ the vector configurations
$\mathbf{G}e_{1},\dots,\mathbf{G}e_{n}\in\mathbb{R}^{d^{\prime}}\qquad\text{and}\qquad\mathbf{H}e_{1},\dots,\mathbf{H}e_{n}\in\mathbb{R}^{d^{\prime\prime}}$
are in linear Gale duality.
We shall give two proofs of Proposition 2.5. The first proof is based on the following general coupling principle.
Lemma 2.6.
Let $X\in\mathbb{R}^{p}$ and $Y\in\mathbb{R}^{q}$ be random vectors, and let $\varphi:\mathbb{R}^{p}\to T$ and $\psi:\mathbb{R}^{q}\to T$ be Borel-measurable maps into a Polish space $T$ (equipped with its Borel $\sigma$-algebra). Assume that $\varphi(X)$ and $\psi(Y)$ have the same distribution on $T$. Then there exists a probability space carrying random vectors $X^{\prime}$ and $Y^{\prime}$ such that
$X^{\prime}\stackrel{{\scriptstyle d}}{{=}}X,\qquad Y^{\prime}\stackrel{{\scriptstyle d}}{{=}}Y,\qquad\text{and}\qquad\varphi(X^{\prime})=\psi(Y^{\prime})\ \text{a.s.}$
Proof.
Let $(\pi_{t}^{X})_{t\in T}$ be a regular conditional distribution of $X$ given $\varphi(X)$, and let $(\pi_{t}^{Y})_{t\in T}$ be a regular conditional distribution of $Y$ given $\psi(Y)$. Existence of these kernels follows since $T$ is Polish; see, e.g., [12, Theorem 10.2.2]. By construction, for every $t\in T$,
(2.4) $\pi_{t}^{X}\bigl(\varphi^{-1}(\{t\})\bigr)=1\qquad\text{and}\qquad\pi_{t}^{Y}\bigl(\psi^{-1}(\{t\})\bigr)=1.$
Now construct a random element $Z$ with values in $T$ such that $Z\stackrel{{\scriptstyle d}}{{=}}\varphi(X)$ (equivalently, $Z\stackrel{{\scriptstyle d}}{{=}}\psi(Y)$). Conditionally on $Z=t$, sample $X^{\prime}$ according to $\pi_{t}^{X}$ and $Y^{\prime}$ according to $\pi_{t}^{Y}$ (for instance, take $X^{\prime}$ and $Y^{\prime}$ conditionally independent given $Z$). Then $X^{\prime}\stackrel{{\scriptstyle d}}{{=}}X$ and $Y^{\prime}\stackrel{{\scriptstyle d}}{{=}}Y$ by the defining property of regular conditional distributions. Moreover, the support properties (2.4) imply $\varphi(X^{\prime})=Z=\psi(Y^{\prime})$ almost surely. ∎
First proof of Proposition 2.5.
We apply Lemma 2.6 with $T:=G(n,d^{\prime\prime})$, $X\stackrel{{\scriptstyle d}}{{=}}\mathbf{G}$ and $Y\stackrel{{\scriptstyle d}}{{=}}\mathbf{H}$, viewed as random elements of $\mathbb{R}^{d^{\prime}\times n}$ and $\mathbb{R}^{d^{\prime\prime}\times n}$, respectively. Define Borel maps
$\varphi:\mathbb{R}^{d^{\prime}\times n}\to T,\quad\varphi(M):=\operatorname*{Ker}(M)\quad\text{ and }\quad\psi:\mathbb{R}^{d^{\prime\prime}\times n}\to T,\quad\psi(M):=\operatorname*{Ker}(M)^{\perp}.$
By (2.2), the random subspaces $\varphi(\mathbf{G})=\operatorname*{Ker}(\mathbf{G})$ and $\psi(\mathbf{H})=\operatorname*{Ker}(\mathbf{H})^{\perp}$ have the same distribution on $T$. Hence Lemma 2.6 yields a coupling of $\mathbf{G}$ and $\mathbf{H}$ on a common probability space such that
(2.5) $\operatorname*{Ker}(\mathbf{G})=\operatorname*{Ker}(\mathbf{H})^{\perp}\quad\text{a.s.}$
The matrices with columns $\mathbf{G}e_{1},\dots,\mathbf{G}e_{n}$ and $\mathbf{H}e_{1},\dots,\mathbf{H}e_{n}$ are precisely $\mathbf{G}$ and $\mathbf{H}$. Since $\mathbf{G}$ and $\mathbf{H}$ are standard Gaussian matrices, they have full row rank almost surely, i.e. $\mathop{\mathrm{rank}}\nolimits(\mathbf{G})=d^{\prime}$ and $\mathop{\mathrm{rank}}\nolimits(\mathbf{H})=d^{\prime\prime}$ a.s. Together with (2.5) this shows that Conditions (2.1) are satisfied, hence the configurations $\mathbf{G}e_{1},\dots,\mathbf{G}e_{n}$ and $\mathbf{H}e_{1},\dots,\mathbf{H}e_{n}$ are in linear Gale duality. ∎
Remark 2.7.
The same argument proves the following more general claim. Consider random matrices $\mathbf{A}:\mathbb{R}^{d^{\prime}+d^{\prime\prime}}\to\mathbb{R}^{d^{\prime}}$ and $\mathbf{B}:\mathbb{R}^{d^{\prime}+d^{\prime\prime}}\to\mathbb{R}^{d^{\prime\prime}}$ such that, both, $\mathbf{A}$ and $\mathbf{B}$ have full rank a.s. and $\operatorname*{Ker}\mathbf{A}$ has the same distribution as $(\operatorname*{Ker}\mathbf{B})^{\bot}$ on the Grassmannian $G(d^{\prime}+d^{\prime\prime},d^{\prime\prime})$. Then, there is a coupling of $\mathbf{A}$ and $\mathbf{B}$ such that $\operatorname*{Ker}\mathbf{A}=(\operatorname*{Ker}\mathbf{B})^{\bot}$ a.s. The proof follows from Lemma 2.6.
Second proof of Proposition 2.5.
In this proof, we give an explicit construction of the coupling $(\mathbf{G},\mathbf{H})$. Let $X_{1},\ldots,X_{d^{\prime\prime}}$ be i.i.d. standard Gaussian points in $\mathbb{R}^{n}=\mathbb{R}^{d^{\prime}+d^{\prime\prime}}$. The linear Blaschke–Petkantschin formula [43, Theorem 7.2.1] states that for every nonnegative Borel function $h:(\mathbb{R}^{n})^{d^{\prime\prime}}\to[0,\infty)$,
$\mathbb{E}h(X_{1},\ldots,X_{d^{\prime\prime}})=\int\displaylimits_{(\mathbb{R}^{n})^{d^{\prime\prime}}}\frac{{\rm e}^{-\frac{1}{2}(\|z_{1}\|^{2}+\ldots+\|z_{d^{\prime\prime}}\|^{2})}}{(2\pi)^{nd^{\prime\prime}/2}}\cdot h(z_{1},\ldots,z_{d^{\prime\prime}})\;\prod_{i=1}^{d^{\prime\prime}}\textup{d}z_{i}\\ =B\int\displaylimits_{G(n,d^{\prime\prime})}\int\displaylimits_{L}\frac{{\rm e}^{-\frac{1}{2}(\|z_{1}\|^{2}+\ldots+\|z_{d^{\prime\prime}}\|^{2})}}{(2\pi)^{nd^{\prime\prime}/2}}\cdot(\nabla_{d^{\prime\prime}}(z_{1},\ldots,z_{d^{\prime\prime}}))^{d^{\prime}}\cdot h(z_{1},\ldots,z_{d^{\prime\prime}})\;\prod_{i=1}^{d^{\prime\prime}}\lambda_{L}(\textup{d}z_{i})\;\nu_{n,d^{\prime\prime}}(\textup{d}L),$
where $\nu_{n,d^{\prime\prime}}$ is the uniform probability distribution on $G(n,d^{\prime\prime})$, $\lambda_{L}$ is the Lebesgue measure on $L\in G(n,d^{\prime\prime})$, $\nabla_{d^{\prime\prime}}(z_{1},\ldots,z_{d^{\prime\prime}})$ denotes the $d^{\prime\prime}$-dimensional volume of the parallelepiped spanned by the vectors $z_{1},\ldots,z_{d^{\prime\prime}}$, and $B$ is a normalizing constant whose value can be found in [43, Eqn. (7.8), p. 271].
To construct the coupling $(\mathbf{G},\mathbf{H})$, we consider the following two-stage random experiment. In a first stage, construct a random Gaussian matrix $\mathbf{G}:\mathbb{R}^{n}\to\mathbb{R}^{d^{\prime}}$. In a second stage, conditionally on $\operatorname*{Ker}\mathbf{G}=L$, where $L\subseteq\mathbb{R}^{n}$ is a $d^{\prime\prime}$-dimensional linear subspace, we construct random points $(Z_{1},\ldots,Z_{d^{\prime\prime}})$ on $L$ whose joint probability density with respect to $(\lambda_{L})^{\otimes d^{\prime\prime}}$ is given by
$f_{L}(z_{1},\ldots,z_{d^{\prime\prime}}):=B\cdot\frac{{\rm e}^{-\frac{1}{2}(\|z_{1}\|^{2}+\ldots+\|z_{d^{\prime\prime}}\|^{2})}}{(2\pi)^{nd^{\prime\prime}/2}}\cdot(\nabla_{d^{\prime\prime}}(z_{1},\ldots,z_{d^{\prime\prime}}))^{d^{\prime}},\qquad z_{1},\ldots,z_{d^{\prime\prime}}\in L.$
(The fact that $f_{L}$ integrates to $1$ for every fixed $L\in G(n,d^{\prime\prime})$ follows from the Blaschke–Petkantschin formula with $h\equiv 1$ together with the fact that the integrand in $\int_{G(n,d^{\prime\prime})}\ldots\nu_{n,d^{\prime\prime}}(\textup{d}L)$ does not depend on $L$.) With probability $1$, the linear span of $Z_{1},\ldots,Z_{d^{\prime\prime}}$ is $L$. Now, for every nonnegative Borel function $h:(\mathbb{R}^{n})^{d^{\prime\prime}}\to[0,\infty)$, the definition of $Z_{1},\ldots,Z_{d^{\prime\prime}}$ yields
$\mathbb{E}h(Z_{1},\ldots,Z_{d^{\prime\prime}})\\ =\int\displaylimits_{G(n,d^{\prime\prime})}\int\displaylimits_{L}B\cdot\frac{{\rm e}^{-\frac{1}{2}(\|z_{1}\|^{2}+\ldots+\|z_{d^{\prime\prime}}\|^{2})}}{(2\pi)^{nd^{\prime\prime}/2}}\cdot(\nabla_{d^{\prime\prime}}(z_{1},\ldots,z_{d^{\prime\prime}}))^{d^{\prime}}\cdot h(z_{1},\ldots,z_{d^{\prime\prime}})\;\prod_{i=1}^{d^{\prime\prime}}\lambda_{L}(\textup{d}z_{i})\;\nu_{n,d^{\prime\prime}}(\textup{d}L).$
So $\mathbb{E}h(Z_{1},\ldots,Z_{d^{\prime\prime}})=\mathbb{E}h(X_{1},\ldots,X_{d^{\prime\prime}})$ for every $h$ and it follows that $(Z_{1},\ldots,Z_{d^{\prime\prime}})$ have the same full joint distribution as $(X_{1},\ldots,X_{d^{\prime\prime}})$, i.e. they are i.i.d. standard Gaussian points in $\mathbb{R}^{n}$. We can now construct $\mathbf{H}:\mathbb{R}^{n}\to\mathbb{R}^{d^{\prime\prime}}$ by declaring that $\mathbf{H}^{\top}$ maps the standard orthonormal basis of $\mathbb{R}^{d^{\prime\prime}}$ to $Z_{1},\ldots,Z_{d^{\prime\prime}}$. Then, $\mathbf{H}$ is a Gaussian matrix and $(\operatorname*{Ker}\mathbf{H})^{\bot}=\operatorname{Im}(\mathbf{H}^{\top})=L=\operatorname*{Ker}\mathbf{G}$. ∎
2.3. Duality coupling for i.i.d. random unit vectors
The next result will be repeatedly used in the subsequent sections.
Theorem 2.8 (Main duality coupling).
Fix some $d,k\in\mathbb{N}$. On a suitable probability space one can construct random vectors $U_{1},\ldots,U_{d+k},V_{1},\ldots,V_{d+k}$ such that
  • (a)
    $U_{1},\ldots,U_{d+k}$ are independent and uniform on the unit sphere $\mathbb{S}^{d-1}$ in $\mathbb{R}^{d}$;
  • (b)
    $V_{1},\ldots,V_{d+k}$ are independent and uniform on the unit sphere $\mathbb{S}^{k-1}$ in $\mathbb{R}^{k}$;
  • (c)
    For every outcome in the underlying probability space and every set $I\subseteq[d+k]$ with $I\leq d-1$ we have
    (2.6) $\mathop{\mathrm{pos}}\nolimits(U_{i}:i\in I)\text{ is a face of }\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d+k})\quad\Longleftrightarrow\quad\mathop{\mathrm{pos}}\nolimits(V_{i}:i\in I^{c})=\mathbb{R}^{k}.$
    Here $\II^{c}:=[d+k]\backslash I$ is the complement of the index set $I$.
Proof.
Let $e_{1},\ldots,e_{d+k}$ be the standard orthonormal basis of $\mathbb{R}^{d+k}$. By Proposition 2.5, we may couple Gaussian projections $\mathbf{G}:\mathbb{R}^{d+k}\to\mathbb{R}^{d}$ and $\mathbf{H}:\mathbb{R}^{d+k}\to\mathbb{R}^{k}$ such that, for almost every realization of $(\mathbf{G},\mathbf{H})$, the vectors $\mathbf{G}e_{1},\ldots,\mathbf{G}e_{d+k}\in\mathbb{R}^{d}$ are in linear Gale duality with the vectors $\mathbf{H}e_{1},\ldots,\mathbf{H}e_{d+k}\in\mathbb{R}^{k}$. Moreover, with probability $1$, both configurations are in general linear position. After restricting to a full-measure event we may assume that these properties hold for all outcomes in the probability space.
We can therefore apply Lemma 2.2 to the Gale-dual configurations $(\mathbf{G}e_{i})_{i\in[d+k]}$ and $(\mathbf{H}e_{i})_{i\in[d+k]}$. It yields that for every set $I\subseteq[d+k]$ with $I\leq d-1$,
(2.7) $\mathop{\mathrm{pos}}\nolimits(\mathbf{G}e_{i}:i\in I)\text{ is a face of }\mathop{\mathrm{pos}}\nolimits(\mathbf{G}e_{i}:i\in[d+k])\quad\Longleftrightarrow\quad\mathop{\mathrm{pos}}\nolimits(\mathbf{H}e_{i}:i\in I^{c})=\mathbb{R}^{k}.$
Now we define $U_{i}:=\mathbf{G}e_{i}/\|\mathbf{G}e_{i}\|$ and $V_{i}:=\mathbf{H}e_{i}/\|\mathbf{H}e_{i}\|$ for all $i=1,\ldots,d+k$. (Note that $\mathbf{G}e_{i}\neq 0$ and $\mathbf{H}e_{i}\neq 0$ by the linear general position property.) Then (a) and (b) are satisfied. Also, for every $M\subseteq[d+k]$ we have
$\mathop{\mathrm{pos}}\nolimits(\mathbf{G}e_{i}:i\in M)=\mathop{\mathrm{pos}}\nolimits(U_{i}:i\in M),\qquad\mathop{\mathrm{pos}}\nolimits(\mathbf{H}e_{i}:i\in M)=\mathop{\mathrm{pos}}\nolimits(V_{i}:i\in M).$
Substituting these identities into (2.7) yields (2.6), completing the proof. ∎
Remark 2.9.
The coupling in Theorem 2.8 is naturally interpreted in the language of oriented matroids [6]. Indeed, for each outcome of the underlying probability space, the vector configurations $(U_{1},\ldots,U_{d+k})\subseteq\mathbb{R}^{d}$ and $(V_{1},\ldots,V_{d+k})\subseteq\mathbb{R}^{k}$ determine oriented matroids on the common ground set $[d+k]$, and the construction in the proof yields that these oriented matroids are dual to one another (in the sense of Gale duality). In the present paper, however, we do not make systematic use of oriented matroid theory: the only consequence of this duality that will be needed is the correspondence stated in part (c) of Theorem 2.8.
3. Two examples
Theorem 2.8 will serve as a key tool throughout the remainder of the paper. We start by illustrating its use with two simple applications.
3.1. Independence of face events
Consider the random cone $\mathcal{W}_{n,d}=\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{n})$, where $U_{1},\ldots,U_{n}$ are independent and uniformly distributed on the unit sphere in $\mathbb{R}^{d}$. For a subset $I\subseteq[n]$ define the cone $F_{I}=\mathop{\mathrm{pos}}\nolimits(U_{i}:i\in I)$ and the random event
$A_{I}:=\{F_{I}\text{ is a face of }\mathcal{W}_{n,d}\}.$
With probability $1$, the vectors $U_{1},\ldots,U_{n}$ are in general linear position (assuming $d\geq 2$). Thus, for $\ell\in\{0,\ldots,d-1\}$, every $\ell$-dimensional face of $\mathcal{W}_{n,d}$ has the form $F_{I}$ for some $I\subseteq[n]$ with $I=\ell$, but not every $F_{I}$ is necessarily a face of $\mathcal{W}_{n,d}$. In particular, the number of $\ell$-dimensional faces of $\mathcal{W}_{n,d}$ can be represented as
$f_{\ell}(\mathcal{W}_{n,d})=\sum_{I\subseteq[n],\#I=\ell}\mathbbm{1}_{A_{I}}.$
Somewhat unexpectedly, many of the face events $A_{I}$ turn out to be independent; the next theorem makes this precise.
Theorem 3.1 (Independence of face events).
Take $d\geq 3$ and $n\geq d+1$. Let $I_{1},\ldots,I_{r}\subseteq[n]$ be such that $I_{s}\cup I_{t}=[n]$ for all $1\leq s\neq t\leq r$ and $I_{s}\leq d-1$ for all $s\in[r]$. Then, the face events $A_{I_{1}},\ldots,A_{I_{r}}$ are independent.
Proof.
We use the coupling given in Theorem 2.8 with k=ndk=n-d. Recall that the random vectors $V_{1},\ldots,V_{n}$ appearing in that theorem are independent and uniformly distributed on the unit sphere in $\mathbb{R}^{n-d}$. For $J\subseteq[n]$ consider the cone $D_{J}=\mathop{\mathrm{pos}}\nolimits(V_{j}:j\in J)\subseteq\mathbb{R}^{n-d}$. By Theorem 2.8,
[…]
Under our assumptions, the index sets $I_{1}^{c},\ldots,I_{r}^{c}$ are disjoint. Hence the random events $\{D_{I_{1}^{c}}=\mathbb{R}^{n-d}\},\ldots,\{D_{I_{r}^{c}}=\mathbb{R}^{n-d}\}$ are independent. ∎
3.2. Expected face counts
As a simple consequence of Theorem 2.8, we recover the following result of Donoho and Tanner [11, Theorem 1.6]. Recall that Wendel’s formula [47] (see also [43, Theorem 8.2.1]) states that
(3.2) $p(n,d)=\mathbb{P}[\mathcal{W}_{n,d}=\mathbb{R}^{d}]=\frac{1}{2^{n-1}}\sum_{\ell=d}^{n-1}\binom{n-1}{\ell}$.
Proposition 3.2 (Expected face numbers).
For every $d\geq 2$, $n\geq d+1$ and $\ell\in\{0,\ldots,d-1\}$, the expected number of $\ell$-dimensional faces of the random cone $\mathcal{W}_{n,d}$ is
$\mathbb{E}f_{\ell}(\mathcal{W}_{n,d})=\binom{n}{\ell}p(n-\ell,n-d)=\frac{1}{2^{n-\ell-1}}\binom{n}{\ell}\sum_{r=n-d}^{n-\ell-1}\binom{n-\ell-1}{r}$.
Also, for every $I\subseteq[n]$ with $I=\ell\in\{0,\ldots,d-1\}$, we have $\mathbb{P}[A_{I}]=p(n-\ell,n-d)$.
Proof.
For every II as above, Formula (3.1) gives
$\mathbb{P}[A_{I}]=\mathbb{P}[\mathop{\mathrm{pos}}\nolimits(V_{i}:i\in I^{c})=\mathbb{R}^{n-d}]=\mathbb{P}[\mathcal{W}_{n-\ell,n-d}=\mathbb{R}^{n-d}]=p(n-\ell,n-d),$
where the last step follows from Wendel’s formula (3.2). Taking the sum over all subsets $I\subseteq[n]$ with $I=\ell$ gives the stated formula for $\mathbb{E}f_{\ell}(\mathcal{W}_{n,d})$. ∎
Remark 3.3.
Taking up a suggestion of Gale, Schneider [41] proposed to study random polytopes $P$ obtained by choosing their (affine) Gale transform at random. (Throughout this remark, we use the term “polytope” to refer to its combinatorial type; only the combinatorial type is canonically determined by the Gale transform.) The model of Schneider [41, p. 643] depends on an even probability measure $\phi$; in the present remark we take $\phi$ to be the uniform distribution on the unit sphere in $\mathbb{R}^{k}$. We put $n=d+k$. Using the notation of Theorem 2.8, Schneider [41] considers a (d1)(d-1)-dimensional random polytope $P$ with $n$ vertices whose (affine) Gale transform is given by $\lambda_{1}V_{1},\ldots,\lambda_{n}V_{n}\in\mathbb{R}^{k}$ (with suitable $\lambda_{j}>0$) conditioned on the event $\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{n})=\mathbb{R}^{k}$. By Example 2.4, this is the same conditioning event as $\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{n})\neq\mathbb{R}^{d}$. It then follows from Theorem 2.8 that the random cone $\mathcal{W}_{n,d}=\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{n})$ conditioned on the event $\{\mathcal{W}_{n,d}\neq\mathbb{R}^{d}\}$, can be identified with the cone over $P$. Conversely, $P$ can be recovered as a bounded slice (cross-section) of this conditioned random cone. Such conditioned cones are called Cover–Efron cones in [20]. Consequently, the random polytopes studied in Schneider [41] have the same distribution as the bounded cross-sections of Cover–Efron cones considered in [20]. (As noted in [41, Remark on p. 648], these two objects in particular have the same expected face numbers.)
4. Moments of random angles
4.1. Duality for moments of random angles
The solid angle of a polyhedral cone with non-empty interior $C\subseteq\mathbb{R}^{d}$ is defined as $\alpha(C)=\nu_{d}(C\cap\mathbb{S}^{d-1})\in[0,1],$ where $\nu_{d}$ is the uniform probability distribution on $\mathbb{S}^{d-1}$.
Let $U_{1},\ldots,U_{d}$ be i.i.d. random unit vectors uniformly distributed on the unit sphere in $\mathbb{R}^{d}$, $d\in\mathbb{N}$. The solid angle of the cone $\mathcal{W}_{d,d}=\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d})$ is a random variable $\alpha_{d}:=\alpha(\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d}))$ taking values in the interval $[0,1/2]$. We are interested in the moments of $\alpha_{d}$, which we denote by
$m(d,k):=\mathbb{E}[\alpha_{d}^{k}],\qquad d,k\in\mathbb{N}.$
Clearly, $m(d,1)=\mathbb{E}\alpha_{d}=2^{-d}$. (Indeed, the cones $\mathop{\mathrm{pos}}\nolimits(\varepsilon_{1}U_{1},\ldots,\varepsilon_{d}U_{d})$, with $\varepsilon_{j}\in\{\pm 1\}$, have disjoint interiors, cover $\mathbb{R}^{d}$ and have the same expected angle.)
Theorem 4.1 (Duality for random solid angles).
For all $d,k\in\mathbb{N}$ we have $m(d,k)=m(k,d)$.
Proof.
Let $U_{1},\ldots,U_{d},U_{d+1},\ldots,U_{d+k}$ be i.i.d. random unit vectors with the uniform distribution on $\mathbb{S}^{d-1}$. We shall compute the probability of the event $U_{d+1},\ldots,U_{d+k}\in\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d})$ in two different ways.
Primal method. Conditioning on $U_{1},\ldots,U_{d}$, we have, by definition of the solid angle,
$\mathbb{P}[U_{d+1},\ldots,U_{d+k}\in\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d})\,|\,U_{1},\ldots,U_{d}]=\left(\alpha(\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d}))\right)^{k}.$
Integrating over $U_{1},\ldots,U_{d}$ we arrive at
$\mathbb{P}[U_{d+1},\ldots,U_{d+k}\in\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d})]=\mathbb{E}[(\alpha(\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d})))^{k}]=m(d,k).$
Dual method. By Theorem 2.8, we may realize $(U_{1},\ldots,U_{d+k})$ together with $(V_{1},\ldots,V_{d+k})$ on a common probability space such that the duality (2.6) holds. In the following we agree to ignore null events. On the event that $\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d})\cap\mathbb{S}^{d-1}$ is a spherical simplex (which holds a.s.), the condition $U_{d+1},\ldots,U_{d+k}\in\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d})$ is equivalent to requiring that all facets of $\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d})$ remain facets of $\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d+k})$. Hence,
$\mathbb{P}\Bigl[U_{d+1},\ldots,U_{d+k}\in\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d})\Bigr]$
$=\mathbb{P}\Bigl[\forall\ell\in[d]:\ \mathop{\mathrm{pos}}\nolimits(U_{i}:i\in[d]\setminus\{\ell\})\text{ is a facet of }\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d+k})\Bigr]$
$=\mathbb{P}\Bigl[\forall\ell\in[d]:\ \mathop{\mathrm{pos}}\nolimits(V_{i}:i\in([d]\setminus\{\ell\})^{c})=\mathbb{R}^{k}\Bigr],$
where we used (2.6). Since $([d]\setminus\{\ell\})^{c}=\{\ell,d+1,\ldots,d+k\}$, this becomes
$\mathbb{P}[U_{d+1},\ldots,U_{d+k}\in\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d})]=\mathbb{P}\Bigl[\forall\ell\in[d]:\ \mathop{\mathrm{pos}}\nolimits(V_{\ell},V_{d+1},\ldots,V_{d+k})=\mathbb{R}^{k}\Bigr].$
Now, we observe that $\mathop{\mathrm{pos}}\nolimits(V_{\ell},V_{d+1},\ldots,V_{d+k})=\mathbb{R}^{k}$ if and only if $V_{\ell}\in-\mathop{\mathrm{Int}}\nolimits\mathop{\mathrm{pos}}\nolimits(V_{d+1},\ldots,V_{d+k})$, ignoring null events. Thus,
$\mathbb{P}[U_{d+1},\ldots,U_{d+k}\in\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d})]$ $=\mathbb{P}[V_{1},\ldots,V_{d}\in-\mathop{\mathrm{pos}}\nolimits(V_{d+1},\ldots,V_{d+k})]$
$=\mathbb{E}[(\alpha(\mathop{\mathrm{pos}}\nolimits(V_{d+1},\ldots,V_{d+k})))^{d}]$
$=m(k,d).$
Comparing the results obtained by both methods, we conclude that $m(d,k)=m(k,d)$. ∎
4.2. Moments of small order
Theorem 4.1 allows us to find explicit formulas for the first three moments of $\alpha_{d}$.
Theorem 4.2 (Moments of random angles).
For every $d\in\mathbb{N}$, the first three moments of the random variable $\alpha_{d}:=\alpha(\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d}))$ are given by
(4.1) $\mathbb{E}[\alpha_{d}]=m(d,1)=m(1,d)=2^{-d},$
(4.2) $\mathbb{E}[\alpha_{d}^{2}]=m(d,2)=m(2,d)=\frac{1}{2^{d}(d+1)},\qquad\mathop{\mathrm{Var}}\nolimits[\alpha_{d}]=\frac{1}{2^{d}(d+1)}-\frac{1}{4^{d}},$
(4.3) $\mathbb{E}[\alpha^{3}_{d}]=m(d,3)=m(3,d)=\frac{1}{(4\pi)^{d}}\mathbb{E}[S^{d}]=\frac{1}{(4\pi)^{d}}\int_{0}^{2\pi}x^{d}f_{S}(x){\rm d}x,$
where $S$ is a random variable with probability density
$f_{S}(x)=-\frac{(x^{2}-4\pi x+3\pi^{2}-6)\cos x-6(x-2\pi)\sin x-2(x^{2}-4\pi x+3\pi^{2}+3)}{16\pi\,\cos^{4}(x/2)},\qquad 0\leq x\leq 2\pi.$
Proof.
To give an alternative proof that $\mathbb{E}[\alpha_{d}]=2^{-d}$, observe that $m(1,d)=2^{-d}$ since $V_{1}$ is uniform on $\{\pm 1\}$ and hence $\mathop{\mathrm{pos}}\nolimits(V_{1})$ is a ray with angle $1/2$.
To prove (4.2), it suffices to compute $m(2,d)$. If $V_{1},V_{2}$ are independent and uniformly distributed on the unit circle in $\mathbb{R}^{2}$, then the random angle $\alpha(\mathop{\mathrm{pos}}\nolimits(V_{1},V_{2}))$ is uniformly distributed on $[0,1/2]$. The $d$-th moment of this random variable is $2^{-d}\cdot 1/(d+1)$, and the claim follows.
Finally, the formula for $m(3,d)$ follows from the following non-trivial fact due to Crofton and Exhumatus [9]: The spherical area $S$ of the random spherical triangle with vertices $V_{1},V_{2},V_{3}$ that are independent and uniform on $\mathbb{S}^{2}$ has probability density $f_{S}$ given as above. Proofs can be found in [38] and [13]; see also [44] and [36]. Then $\alpha(\mathop{\mathrm{pos}}\nolimits(V_{1},V_{2},V_{3}))=S/(4\pi)$ and $m(3,d)=\mathbb{E}[S^{d}/(4\pi)^{d}]$. ∎
Remark 4.3 (The third moment).
Explicit formulas for $\mathbb{E}[S^{d}]$ with $d\leq 10$ are given in [44]. (In the formula for $\mathbb{E}[S^{4}]$, a factor of $108$ is missing in front of $\zeta(3)$.) These formulas give the trivial values $\mathbb{E}[\alpha_{1}^{3}]=\frac{1}{8}$, $\mathbb{E}[\alpha_{2}^{3}]=\frac{1}{32}$ (since $\alpha_{2}$ is uniform on $[0,1/2]$) as well as
$\mathbb{E}[\alpha_{3}^{3}]=\frac{3}{128}\;-\;\frac{3\log 2}{16\pi^{2}},\qquad\mathbb{E}[\alpha_{4}^{3}]=\frac{\pi^{4}-6\pi^{2}\log 2-27\,\zeta(3)}{64\,\pi^{4}},\qquad\mathbb{E}[\alpha_{5}^{3}]=\frac{5}{512}\;-\;\frac{15\log 2}{128\pi^{2}}.$
Remark 4.4 (Asymptotics of the third moments).
Since $f_{S}(2\pi)=\frac{12-\pi^{2}}{16\pi}>0$ and $f_{S}(x)$ is continuous at $x=2\pi$, the standard Laplace asymptotics gives
$\mathbb{E}[\alpha^{3}_{d}]=\frac{1}{(4\pi)^{d}}\mathbb{E}[S^{d}]\sim\frac{1}{(4\pi)^{d}}\frac{f_{S}(2\pi)\,(2\pi)^{d+1}}{d+1}\sim\frac{12-\pi^{2}}{8}\frac{1}{2^{d}d},\qquad d\to\infty.$
Remark 4.5.
In this section, we considered the cone $\mathcal{W}_{n,d}$ with $n=d$ only. Our methods do not extend to $n>d$. For example, it remains open to compute the variance of $\alpha(\mathcal{W}_{d+1,d})$.
4.3. Conjectures on high-dimensional random angles
It seems difficult to compute $m(d,k)=m(k,d)$ with $k,d\geq 4$. Here we state a conjecture on the asymptotics of $m(d,k)$ as $d\to\infty$ with fixed $k\geq 2$. Let $V_{1},\ldots,V_{k}$ be independent and uniform on $\mathbb{S}^{k-1}$. The maximal possible value of the random variable $\alpha(\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{k}))$ is $1/2$. An analysis of configurations of $k$ unit vectors in $\mathbb{R}^{k}$ close to being linearly dependent suggests that the density of $\alpha(\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{k}))$ at $1/2$ should be positive, which (arguing as in Remark 4.4) leads to the following conjecture.
Conjecture 4.6 (Asymptotics for higher moments).
For every fixed $k\in\{2,3,\ldots\}$, we have $m(d,k)\sim c_{k}/(2^{d}\cdot d)$ as $d\to\infty$, where $c_{k}>0$ is a constant. For example, $c_{2}=1$ and $c_{3}=(12-\pi^{2})/8$. (Note that the case $k=1$ is different since $m(d,1)=2^{-d}$.)
This behavior of moments suggests that, for large $d$, the random variable $\alpha_{d}=\alpha(\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d}))$ can be approximated by a mixture of a random “background” of size $2^{-d}$ and rare random “spikes” with size of order $1$ and probability of order $1/(d\cdot 2^{d})$. More precisely, let $A_{d}$ be an event with $\mathbb{P}[A_{d}]=1/(d\cdot 2^{d})$ and let $X,Y\geq 0$ be random variables with finite moments. Then, the random variables $Z_{d}:=(1-\mathbbm{1}_{A_{d}})2^{-d}X+\mathbbm{1}_{A_{d}}Y$ have moments whose large-$d$ behavior is similar to that of $\alpha_{d}$, namely $\mathbb{E}\,[Z_{d}]\sim 2^{-d}\,\mathbb{E}\,[X]$ as $d\to\infty$ and $\mathbb{E}\,[Z_{d}^{k}]\sim\mathbb{E}\,[Y^{k}]/(d\cdot 2^{d})$ for $k\geq 2$. This suggests the following conjectures.
Conjecture 4.7 (Weak limit for random high-dimensional angle).
As $d\to\infty$, the random variables $2^{d}\alpha_{d}$ converge in distribution to some random variable $X\geq 0$ with $\mathbb{E}\,[X]=1$.
Conjecture 4.8 (Rare spikes for high-dimensional random angle).
The sequence of measures $(\mu_{d})_{d\in\mathbb{N}}$ with $\mu_{d}(A):=2^{d}d\cdot\mathbb{P}[\alpha_{d}\in A]$, $A\subseteq(0,\infty)$, converges weakly to a non-trivial limit.
5. A spherical Sylvester problem
Take some $d\geq 2$, $k\in\mathbb{N}$, and let $U_{1},\ldots,U_{d+k}$ be i.i.d. random unit vectors uniformly distributed on the unit sphere in $\mathbb{R}^{d}$. We consider the random spherical polytope
$\mathcal{P}_{d+k,d}:=\mathcal{W}_{d+k,d}\cap\mathbb{S}^{d-1}=\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d+k})\cap\mathbb{S}^{d-1}.$
Wendel’s formula (1.1) gives the probability that $\mathcal{P}_{d+k,d}$ is the whole sphere.
Proposition 5.1 (Probability of spherical simplex).
For $d\geq 2$, $k\in\mathbb{N}$, the probability that $\mathcal{P}_{d+k,d}$ is a spherical simplex (with $d$ vertices) equals
$\binom{d+k}{d}m(d,k)=\binom{d+k}{k}m(k,d).$
Proof.
The event $\{\mathcal{P}_{d+k,d}\text{ is a spherical simplex}\}$ occurs if and only if some (necessarily unique, for $d\geq 2$) set of $k$ vectors $U_{i_{1}},\ldots,U_{i_{k}}$ is contained in the positive hull of the remaining $d$ vectors. By exchangeability, it follows that
$\mathbb{P}[\mathcal{P}_{d+k,d}\text{ is a spherical simplex}]$ $=\binom{d+k}{d}\cdot\mathbb{P}[U_{d+1},\ldots,U_{d+k}\in\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d})]$
$=\binom{d+k}{d}\cdot\mathbb{E}[(\alpha(\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d})))^{k}]=\binom{d+k}{d}m(d,k),$
and the proof is complete since $m(d,k)=m(k,d)$ by Theorem 4.1. ∎
Example 5.2.
Taking $k=1,2,3$ and recalling Theorem 4.2 yields the following results:
(5.1) $\mathbb{P}[\mathcal{P}_{d+1,d}\text{ is a spherical simplex}]=\frac{d+1}{2^{d}},$
(5.2) $\mathbb{P}[\mathcal{P}_{d+2,d}\text{ is a spherical simplex}]=\frac{d+2}{2^{d+1}},$
(5.3) $\mathbb{P}[\mathcal{P}_{d+3,d}\text{ is a spherical simplex}]=\frac{1}{(4\pi)^{d}}\binom{d+3}{3}\int_{0}^{2\pi}x^{d}f_{S}(x){\rm d}x,$
where $f_{S}(x)$ is as in Theorem 4.2. Equation (5.1) recovers a formula obtained in [31, Lemma 4.4]; see also [30, Theorem 3.6], [32, Corollary 8.1], [23, Section 4.2]. Note that (5.1) admits a simple proof: $\mathcal{P}_{d+1,d}$ is a spherical simplex if one of the vectors ($d+1$ choices) is inside the positive hull of the remaining vectors, and the probability that $U_{d+1}\in\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d})$ is the expected angle of $\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d})$, which is equal to $2^{-d}$ for symmetry reasons (see [31, Section 4] for details). On the other hand, Equations (5.2) and (5.3) seem to be new. For example, (5.3) combined with Remark 4.3 gives
$\mathbb{P}[\mathcal{P}_{6,3}\text{ is a spherical simplex}]=\frac{15}{32}-\frac{15\log 2}{4\pi^{2}}\approx 0.205385,$
$\mathbb{P}[\mathcal{P}_{7,4}\text{ is a spherical simplex}]=\frac{35}{64}-\frac{105\log 2}{32\pi^{2}}-\frac{945\,\zeta(3)}{64\pi^{4}}\approx 0.134219,$
$\mathbb{P}[\mathcal{P}_{8,5}\text{ is a spherical simplex}]=\frac{35}{64}\;-\;\frac{105\,\log 2}{16\pi^{2}}\approx 0.085987.$
Next we consider the case $n=d+2$ in more detail. The number of vertices of the spherical polytope $\mathcal{P}_{d+2,d}$ can take values $d$ (if $\mathcal{P}_{d+2,d}$ is a spherical simplex), $d+1$, $d+2$ and 0 (equivalently, $\mathcal{P}_{d+2,d}=\mathbb{S}^{d-1}$). In the next theorem we compute the probabilities of these possibilities.
Theorem 5.3 ($d+2$ points on $\mathbb{S}^{d-1}$).
For all $d\geq 2$ we have
(5.4) $\mathbb{P}[\mathcal{P}_{d+2,d}=\mathbb{S}^{d-1}]=\frac{d+2}{2^{d+1}},$
(5.5) $\mathbb{P}[\mathcal{P}_{d+2,d}\text{ is a spherical simplex}]=\frac{d+2}{2^{d+1}},$
(5.6) $\mathbb{P}[\mathcal{P}_{d+2,d}\text{ has }d+1\text{ vertices}]=\frac{(d-2)(d+2)}{2^{d+1}},$
(5.7) $\mathbb{P}[\mathcal{P}_{d+2,d}\text{ has }d+2\text{ vertices}]=1-\frac{d(d+2)}{2^{d+1}}.$
Proof.
By Wendel’s formula (3.2), $\mathbb{P}[\mathcal{P}_{d+2,d}=\mathbb{S}^{d-1}]=p(d+2,d)=(d+2)/2^{d+1}$, which proves (5.4). We already proved (5.5) in Example 5.2. It remains to prove (5.6) and (5.7). By Proposition 3.2 with $\ell=1$ (taking into account that $f_{0}(\mathcal{P}_{d+2,d})=f_{1}(\mathcal{W}_{d+2,d})$),
$\mathbb{E}[d+2-f_{0}(\mathcal{P}_{d+2,d})]=(d+2)(1-p(d+1,2))=(d+2)p(d+1,d-1)=\frac{(d+2)(2d+2)}{2^{d+1}},$
where we again used Wendel’s formula (3.2). On the other hand, by definition of expectation and (5.5),
[…]
Comparing both results gives (5.6). To prove (5.7), observe that the complement of $\{\mathcal{P}_{d+2,d}\text{ has }d+2\text{ vertices}\}$ is the disjoint union of the events appearing in (5.4), (5.5), (5.6). ∎
Example 5.4 ($5$ points on $\mathbb{S}^{2}$).
Taking $d=3$ in Theorem 5.3 gives: The probability that the spherical convex hull of five random points on $\mathbb{S}^{2}$, chosen uniformly and independently, is a spherical triangle, quadrilateral, or pentagon is $5/16,5/16,1/16$, respectively. (Also, with probability $5/16$, their positive hull is the whole of $\mathbb{R}^{3}$). Elementary proofs can be found in [30, Theorem 3.6], [32, Exercises 8.1, 8.2 on p. 170], [44].
Let us now consider the regime where $d\to\infty$ while $k\geq 2$ is fixed. Conjecture 4.6, together with Proposition 5.1, would imply that the probability that $\mathcal{P}_{d+k,d}$ is a spherical simplex is asymptotically equivalent to $(c_{k}/k!)\,d^{k-1}2^{-d}$ for some constant $c_{k}>0$. Section 4.3 suggests the following mechanism by which this rare event may occur: a subset of $d$ vectors $U_{i_{1}},\ldots,U_{i_{d}}$ spans a cone with solid angle of constant order, and the remaining $k$ vectors fall inside this cone. This leads to the following conjecture.
Conjecture 5.5.
Fix $k\geq 2$. The conditional distribution of $\alpha(\mathcal{W}_{d+k,d})$, given that $\mathcal{W}_{d+k,d}\cap\mathbb{S}^{d-1}$ is a spherical simplex, converges weakly (as $d\to\infty$) to a limit law not concentrated at 0.
Let us record one further consequence of Theorem 2.8.
Proposition 5.6 (Probability of $\ell$-neighborliness).
Let $d,k\in\mathbb{N}$. Let $U_{1},\ldots,U_{d+k}$ be independent and uniform on $\mathbb{S}^{d-1}$ and let $V_{1},\ldots,V_{d+k}$ be independent and uniform on $\mathbb{S}^{k-1}$. For every $\ell\in\{0,\ldots,d-1\}$, the following events have the same probability:
  • (a)
    $f_{\ell}(\mathcal{W}_{d+k,d})=\binom{d+k}{\ell}$, that is, for every subset $I\subseteq[d+k]$ with $I=\ell$, the cone $\mathop{\mathrm{pos}}\nolimits(U_{i}:i\in I)$ is an $\ell$-dimensional face of $\mathcal{W}_{d+k,d}=\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{d+k})$.
  • (b)
    For every $J\subseteq[d+k]$ with $J=d+k-\ell$, one has $\bigcap_{j\in J}\mathrm{HS}_{j}=\varnothing$, where $\mathrm{HS}_{j}:=\{x\in\mathbb{S}^{k-1}:\langle x,V_{j}\rangle\leq 0\}$, $j=1,\ldots,d+k$. Equivalently, the family of half-spheres $(\mathrm{HS}_{j})_{j=1}^{d+k}$ covers every point of $\mathbb{S}^{k-1}$ at most $d+k-\ell-1$ times.
  • (c)
    The family $(\mathrm{HS}_{j})_{j=1}^{d+k}$ covers every point of $\mathbb{S}^{k-1}$ at least $\ell+1$ times.
Proof.
By Theorem 2.8 (c), the event in (a) has the same probability as the event that for every $J\subseteq[d+k]$ with $J=d+k-\ell$ one has $\mathop{\mathrm{pos}}\nolimits(V_{j}:j\in J)=\mathbb{R}^{k}$. By polarity, $\mathop{\mathrm{pos}}\nolimits(V_{j}:j\in J)\neq\mathbb{R}^{k}$ holds if and only if there exists $x\in\mathbb{S}^{k-1}$ with $\langle x,V_{j}\rangle\leq 0$ for all $j\in J$, that is, if and only if $\bigcap_{j\in J}\mathrm{HS}_{j}\neq\varnothing$. Hence $\mathop{\mathrm{pos}}\nolimits(V_{j}:j\in J)=\mathbb{R}^{k}$ is equivalent to $\bigcap_{j\in J}\mathrm{HS}_{j}=\varnothing$, which is precisely (b).
To see that the events in (b) and (c) have equal probabilities, observe that
Event in (b) $\;\Longleftrightarrow\;\forall x\in\mathbb{S}^{k-1}:\#\{j\in[d+k]:x\in\mathrm{HS}_{j}\}\leq d+k-\ell-1$
$\;\Longleftrightarrow\;\forall x\in\mathbb{S}^{k-1}:\#\{j\in[d+k]:x\in\mathbb{S}^{k-1}\backslash\mathrm{HS}_{j}\}\geq\ell+1.$
Since $\mathbb{S}^{k-1}\backslash\mathrm{HS}_{j}=\{z\in\mathbb{S}^{k-1}:\langle z,-V_{j}\rangle<0\}$ and since $(V_{1},\ldots,V_{d+k})$ and $(-V_{1},\ldots,-V_{d+k})$ have the same distribution, the probability of the latter event equals the probability of (c). ∎
6. Limit theorems for face counts
6.1. Statement of the limit theorem
Recall that we consider the random polyhedral cone $\mathcal{W}_{n,d}=\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{n})$, where $U_{1},\ldots,U_{n}$ are independent and uniformly distributed on the unit sphere in $\mathbb{R}^{d}$. Not much seems to be known about limit distributions of $f_{\ell}(\mathcal{W}_{n,d})$ as $n,d,\ell\to\infty$. We now describe the asymptotic fluctuations of the (dq)(d-q)-dimensional face count of $\mathcal{W}_{d+k,d}$ in the regime $d\to\infty$ with $k\in\mathbb{N}$ and $q\in\mathbb{N}$ fixed. Recall from Proposition 3.2 the explicit formula
$\mathbb{E}f_{d-q}(\mathcal{W}_{d+k,d})=\binom{d+k}{d-q}p(k+q,k)=\frac{1}{2^{k+q-1}}\binom{d+k}{d-q}\sum_{r=k}^{k+q-1}\binom{k+q-1}{r},\qquad d\geq q.$
Theorem 6.1 (Distributional limit for face counts).
Let $k\in\mathbb{N}$ and $q\in\mathbb{N}$ be fixed. Then, as $d\to\infty$, the number of (dq)(d-q)-dimensional faces of $\mathcal{W}_{d+k,d}$ satisfies
(6.1) $d\cdot\left(\frac{f_{d-q}(\mathcal{W}_{d+k,d})}{\binom{d+k}{d-q}}-p(k+q,k)\right)\;{\overset{w}{\underset{d\to\infty}{\longrightarrow}}}\;\frac{\binom{k+q}{2}\binom{k+q-2}{k-1}}{2^{k+q-1}}\,(1-Q_{k}),$
where $Q_{k}>0$ is a random variable admitting the distributional representation $Q_{1}\stackrel{{\scriptstyle d}}{{=}}\operatorname{Gamma}(\frac{1}{2},\frac{1}{2})$ (for $k=1$) and
(6.2) $Q_{k}\stackrel{{\scriptstyle d}}{{=}}\sum_{n=1,3,5,\ldots}\left(\frac{\Gamma\!\left(\frac{k}{2}\right)\Gamma\!\left(\frac{n}{2}\right)}{\pi\,\Gamma\!\left(\frac{n+k}{2}\right)}\right)^{2}\operatorname{Gamma}\left(\frac{d_{n,k}}{2},\frac{1}{2}\right),\qquad\text{ for }k\geq 2.$
Here $d_{n,k}:=\binom{n+k-1}{n}-\binom{n+k-3}{n-2}$ with $k\geq 2$, $n\in\mathbb{N}$ is the dimension of the space of degree $n$ spherical harmonics on $\mathbb{S}^{k-1}$ (see Section 6.6), and $\operatorname{Gamma}(\alpha,\lambda)$ denotes a Gamma-variable with shape parameter $\alpha>0$, rate parameter $\lambda>0$, and all such variables are independent. The expectation and variance of $Q_{k}$ are given by
(6.3) $\mathbb{E}Q_{k}=1,\qquad\mathop{\mathrm{Var}}\nolimits Q_{k}=\frac{4}{\pi^{2}}\,\psi_{1}\left(\frac{k}{2}\right),$
where $\psi_{1}$ is the trigamma function, defined for $x>0$ by $\psi_{1}(x):=\frac{\textup{d}^{2}}{\textup{d}x^{2}}\log\Gamma(x)=\sum_{j=0}^{\infty}\frac{1}{(j+x)^{2}}.$ Finally, as $d\to\infty$ we have
(6.4) $\mathop{\mathrm{Var}}\nolimits\left(\frac{f_{d-q}(\mathcal{W}_{d+k,d})}{\binom{d+k}{d-q}}\right)\sim\frac{\binom{k+q}{2}^{2}\binom{k+q-2}{k-1}^{2}}{2^{2k+2q-2}}\frac{\mathop{\mathrm{Var}}\nolimits Q_{k}}{d^{2}}=\frac{\binom{k+q}{2}^{2}\binom{k+q-2}{k-1}^{2}\psi_{1}\left(\frac{k}{2}\right)}{2^{2k+2q-4}\pi^{2}\cdot d^{2}}.$
Example 6.2 ($k=2$).
For $k=2$, we have $d_{n,2}=2$ for $n\geq 1$. The distributional representation (6.2) simplifies to
$Q_{2}\;\stackrel{{\scriptstyle d}}{{=}}\;\frac{8}{\pi^{2}}\sum_{n=1,3,5,\dots}\frac{E_{n}}{n^{2}},$
where $E_{1},E_{3},E_{5},\dots$ are i.i.d. exponential with unit mean. It follows that the Laplace transform of $Q_{2}$ is given by
$\mathbb{E}\!\left[{\rm e}^{-sQ_{2}}\right]=\prod_{n=1,3,5,\ldots}\mathbb{E}\!\left[\exp\!\left(-s\frac{8}{\pi^{2}}\frac{E_{n}}{n^{2}}\right)\right]=\prod_{n=1,3,5,\ldots}\frac{1}{1+\frac{8s}{\pi^{2}n^{2}}}=\frac{1}{\cosh\!\big(\sqrt{2s}\big)},\qquad s\geq 0.$
This distribution is well known and its properties have been reviewed in [5]. In particular, the following interpretation is known. Let $(B_{t})_{t\geq 0}$ be standard Brownian motion with $B_{0}=0$, and define the first exit time from (1,1)(-1,1) by $\tau_{1}\;:=\;\inf\{t\geq 0:\ |B_{t}|=1\}$. Then $\tau_{1}$ has the same law as $Q_{2}$.
Let us record a convex-dual analogue of Theorem 6.1. Let $U_{1},\ldots,U_{n}$ be independent random vectors, each uniformly distributed on $\mathbb{S}^{d-1}$, where $n>d$. The hyperplanes $U_{1}^{\perp},\ldots,U_{n}^{\perp}$ induce a conical tessellation of $\mathbb{R}^{d}$ into polyhedral cones. The number of cones is almost surely constant and equals $C(n,d)=2\sum_{\ell=0}^{d-1}\binom{n-1}{\ell}$, by the Steiner–Schläfli formula; see [43, Lemma 8.2.1]. Choose one of these cones uniformly at random and denote it by $\mathcal{S}_{n,d}$. The random cone $\mathcal{S}_{n,d}$ is called the random Schläfli cone; see [20, Definition 3.2].
Corollary 6.3 (Distributional limit for face counts of Schläfli cones).
Fix $k\in\mathbb{N}$ and $q\in\mathbb{N}$. Then, as $d\to\infty$, the number of $q$-dimensional faces of $\mathcal{S}_{d+k,d}$ satisfies
(6.5) $d\cdot\left(\frac{f_{q}(\mathcal{S}_{d+k,d})}{\binom{d+k}{\,d-q\,}}-p(k+q,k)\right)\;{\overset{w}{\underset{d\to\infty}{\longrightarrow}}}\;\frac{\binom{k+q}{2}\binom{k+q-2}{k-1}}{2^{k+q-1}}\,(1-Q_{k}),$
where $Q_{k}>0$ is the same random variable as in Theorem 6.1.
Proof.
Let $\mathcal{W}_{d+k,d}^{\circ}$ denote the convex dual of $\mathcal{W}_{d+k,d}$. Then $f_{q}(\mathcal{W}_{d+k,d}^{\circ})=f_{d-q}(\mathcal{W}_{d+k,d})$, and Theorem 6.1 can be rewritten as
(6.6) $d\cdot\left(\frac{f_{q}(\mathcal{W}_{d+k,d}^{\circ})}{\binom{d+k}{\,d-q\,}}-p(k+q,k)\right)\;{\overset{w}{\underset{d\to\infty}{\longrightarrow}}}\;\frac{\binom{k+q}{2}\binom{k+q-2}{k-1}}{2^{k+q-1}}\,(1-Q_{k}).$
By [20, Theorem 3.1], $\mathcal{S}_{d+k,d}$ has the same distribution as $\mathcal{W}_{d+k,d}^{\circ}$ conditioned on the event $\{\mathcal{W}_{d+k,d}\neq\mathbb{R}^{d}\}$. Moreover, Wendel’s formula (1.1) yields
$\mathbb{P}[\mathcal{W}_{d+k,d}\neq\mathbb{R}^{d}]=1-p(d+k,d)=1-\frac{1}{2^{d+k-1}}\sum_{\ell=d}^{d+k-1}\binom{d+k-1}{\ell}\to 1\qquad\text{as }d\to\infty.$
Hence conditioning on $\{\mathcal{W}_{d+k,d}\neq\mathbb{R}^{d}\}$ becomes asymptotically negligible, and we may replace $f_{q}(\mathcal{W}_{d+k,d}^{\circ})$ by $f_{q}(\mathcal{S}_{d+k,d})$ in (6.6) without affecting convergence in distribution. ∎
6.2. Background from the theory of $U$-statistics
The proof of Theorem 6.1 relies on the theory of $U$-statistics. In this section, we introduce the necessary notation and recall several classical results on $U$-statistics; see the books of Lee [29], Korolyuk and Borovskikh [27], and Serfling [45, Chapter 5] for further background.
Let $X_{1},X_{2},\ldots$ be i.i.d. random elements in a measurable space $(\mathcal{X},\mathcal{A})$ with common law $\mu$. Let $h:\mathcal{X}^{m}\to\mathbb{R}$ be a kernel of order $m\in\mathbb{N}$. Our standing assumption is that $h$ is measurable, symmetric (invariant under permutations of its arguments), and
$\mathbb{E}\bigl[h(X_{1},\ldots,X_{m})^{2}\bigr]<\infty.$
For $n\geq m$, the $U$-statistic associated with $h$ is
$\mathbb{U}_{n}:=\binom{n}{m}^{-1}\sum_{1\leq i_{1}<\cdots<i_{m}\leq n}h\!\left(X_{i_{1}},\ldots,X_{i_{m}}\right).$
Put
$g_{0}:=\mathbb{E}h(X_{1},\ldots,X_{m})\qquad(\text{so }\mathbb{E}\mathbb{U}_{n}=g_{0}\text{ for all }n\geq m).$
For $c\in\{1,\ldots,m\}$ define the conditional expectation kernels $g_{c}:\mathcal{X}^{c}\to\mathbb{R}$ by
$g_{c}(x_{1},\ldots,x_{c}):=\mathbb{E}\!\left[h(x_{1},\ldots,x_{c},X_{c+1},\ldots,X_{m})\right],\qquad(x_{1},\ldots,x_{c})\in\mathcal{X}^{c},$
where $X_{c+1},\ldots,X_{m}$ are i.i.d. with law $\mu$. Note that $\mathbb{E}g_{c}(X_{1},\ldots,X_{c})=g_{0}$ and define
$\widetilde{g}_{c}(x_{1},\ldots,x_{c})=g_{c}(x_{1},\ldots,x_{c})-g_{0}.$
Since $h$ is symmetric, each $g_{c}$ (and $\widetilde{g}_{c}$) is symmetric in its arguments. For $c\in\{1,\ldots,m\}$, define
$\zeta_{c}:=\mathop{\mathrm{Var}}\nolimits g_{c}(X_{1},\ldots,X_{c})=\mathbb{E}\bigl[\widetilde{g}_{c}(X_{1},\ldots,X_{c})^{2}\bigr].$
For $n\geq m$, the variance of $\mathbb{U}_{n}$ is given by the exact formula (see [29, Chapter 1, § 1.3, Theorem 3] or [45, Section 5.2.1, Lemma A])
(6.7) $\mathop{\mathrm{Var}}\nolimits\mathbb{U}_{n}=\sum_{c=1}^{m}\frac{\binom{m}{c}\binom{n-m}{\,m-c\,}}{\binom{n}{m}}\;\zeta_{c}.$
The kernel $h$ (and hence $\mathbb{U}_{n}$) is said to be degenerate of order $c\in\{1,\ldots,m\}$ if
$\zeta_{1}=\cdots=\zeta_{c-1}=0\qquad\text{and}\qquad\zeta_{c}>0.$
Equivalently, $g_{r}(x_{1},\ldots,x_{r})=g_{0}$ for all $r\in\{1,\ldots,c-1\}$ ($\mu^{r}$-a.e.), but $g_{c}(X_{1},\ldots,X_{c})$ is not a.s. constant. If $h$ is degenerate of order $c$, Equation (6.7) implies (see, e.g., [45, Section 5.3.4])
(6.8) $\mathop{\mathrm{Var}}\nolimits\mathbb{U}_{n}=\frac{\binom{m}{c}\binom{n-m}{m-c}}{\binom{n}{m}}\;\zeta_{c}\;+\;o(n^{-c})\sim\frac{\binom{m}{c}^{2}\,c!}{n^{c}}\;\zeta_{c},\qquad n\to\infty.$
The next classical theorem characterizes the weak limit of the $U$-statisic $\mathbb{U}_{n}$ under degeneracy of order $2$; see [29, Section 3.2.2], [27, Chapter 4], [40, p. 168] or [45, Section 5.5.2, p. 194].
Theorem 6.4 (Non-Gaussian weak limit under degeneracy of order $2$).
Assume $h:\mathcal{X}^{m}\to\mathbb{R}$ is degenerate of order $2$ (equivalently, $\zeta_{1}=0$ and $\zeta_{2}>0$). Define the (self-adjoint, Hilbert–Schmidt, hence compact) operator $T:L^{2}(\mathcal{X},\mu)\to L^{2}(\mathcal{X},\mu)$ by
$(Tf)(x):=\int_{\mathcal{X}}\widetilde{g}_{2}(x,y)\,f(y)\,\mu(\textup{d}y).$
Let $\lambda_{1},\lambda_{2},\ldots$ be the (real) non-zero eigenvalues of $T$, listed with multiplicities, and let $\xi_{1},\xi_{2},\ldots$ be i.i.d. standard normal random variables. Then
$n\,(\mathbb{U}_{n}-\mathbb{E}\mathbb{U}_{n})\overset{w}{\underset{n\to\infty}{\longrightarrow}}\binom{m}{2}\,\sum_{j=1}^{\infty}\lambda_{j}\,(\xi_{j}^{2}-1).$
The series $\sum_{j=1}^{\infty}\lambda_{j}\,(\xi_{j}^{2}-1)$ converges in $L^{2}$ and a.s. (since $\sum_{j=1}^{\infty}\lambda_{j}^{2}<\infty$ by the Hilbert–Schmidt property of $T$ and $\mathbb{E}[\xi_{j}^{4}]<\infty$) and hence defines a proper random limit.
6.3. Proof of Theorem 6.1: Representation of face count as $U$-statistic
Sections 6.36.8 are devoted to the proof of Theorem 6.1.
The first step is to interpret $f_{\ell}(\mathcal{W}_{n,d})/\binom{n}{\ell}$ as a $U$-statistic, which becomes possible after passing to the Gale dual. Consider the random cone $\mathcal{W}_{n,d}$ with $n>d\geq 2$ and let $\ell\in\{0,\ldots,d-1\}$. With probability $1$, the vectors $U_{1},\ldots,U_{n}$ spanning this cone are in general linear position. Thus, every $\ell$-dimensional face of $\mathcal{W}_{n,d}$ has the form $\mathop{\mathrm{pos}}\nolimits(U_{i}:i\in I)$ for some unique set $I\subseteq[n]$ with $I=\ell$ and hence
$f_{\ell}(\mathcal{W}_{n,d})=\sum_{\begin{subarray}{c}I\subseteq[n]\\ \#I=\ell\end{subarray}}\mathbbm{1}_{\{\mathop{\mathrm{pos}}\nolimits(U_{i}:i\in I)\text{ is a face of }\mathcal{W}_{n,d}\}}.$
We now use the coupling given in Theorem 2.8 (with k=ndk=n-d). Recall that the random vectors $V_{1},\ldots,V_{n}$ appearing in that theorem are independent and uniformly distributed on the unit sphere $\mathbb{S}^{n-d-1}\subseteq\mathbb{R}^{n-d}$. By Theorem 2.8, we have the following equality of random events:
$\{\mathop{\mathrm{pos}}\nolimits(U_{i}:i\in I)\text{ is a face of }\mathcal{W}_{n,d}\}=\{\mathop{\mathrm{pos}}\nolimits(V_{j}:j\in I^{c})=\mathbb{R}^{n-d}\},$
for all $I\subseteq[n]$ with $I\leq d-1$. Putting $J:=I^{c}$ and dividing by $\binom{n}{\ell}$ we arrive at the representation
(6.9) $\frac{f_{\ell}(\mathcal{W}_{n,d})}{\binom{n}{\ell}}=\binom{n}{n-\ell}^{-1}\sum_{\begin{subarray}{c}J\subseteq[n]\\ \#J=n-\ell\end{subarray}}\mathbbm{1}_{\{\mathop{\mathrm{pos}}\nolimits(V_{j}:j\in J)=\mathbb{R}^{n-d}\}},\qquad\ell\in\{0,\ldots,d-1\}.$
The right-hand side is a $U$-statistic with the sample space $\mathcal{X}=\mathbb{S}^{n-d-1}$ (endowed with the uniform distribution $\nu_{n-d}$) and the kernel $h:\mathcal{X}^{n-\ell}\to\{0,1\}$ given by
$h(v_{1},\ldots,v_{n-\ell})=\mathbbm{1}_{\{\mathop{\mathrm{pos}}\nolimits(v_{1},\ldots,v_{n-\ell})=\mathbb{R}^{n-d}\}},\qquad v_{1},\ldots,v_{n-\ell}\in\mathbb{S}^{n-d-1}.$
Recall that in Theorem 6.1 we consider $f_{\ell}(\mathcal{W}_{n,d})$, where the number of points is $n=d+k$, the dimension of faces is $\ell=d-q$, and the parameters $k\in\mathbb{N}$ and $q\in\mathbb{N}$ are fixed. Let us summarize our setting:
  • The sample space is $\mathcal{X}=\mathbb{S}^{k-1}$, endowed with the uniform distribution $\nu_{k}$.
  • $V_{1},\ldots,V_{n}$, where $n=d+k$, are i.i.d. random vectors uniformly distributed on $\mathbb{S}^{k-1}$.
  • The order of the $U$-statistic is $m:=n-\ell=k+q$ and the kernel $h:\mathcal{X}^{m}\to\{0,1\}$ is
    $h(v_{1},\ldots,v_{m})=\mathbbm{1}_{\{\mathop{\mathrm{pos}}\nolimits(v_{1},\ldots,v_{m})=\mathbb{R}^{k}\}},\qquad v_{1},\ldots,v_{m}\in\mathbb{S}^{k-1}.$
By (6.9) we have the identification
(6.10) $\mathbb{U}_{n}=\binom{d+k}{m}^{-1}\sum_{1\leq i_{1}<\cdots<i_{m}\leq d+k}\mathbbm{1}_{\{\mathop{\mathrm{pos}}\nolimits(V_{i_{1}},\ldots,V_{i_{m}})=\mathbb{R}^{k}\}}=\frac{f_{d-q}(\mathcal{W}_{d+k,d})}{\binom{d+k}{d-q}}$.
Our aim is apply Theorem 6.4 to $\mathbb{U}_{n}$. In the following Sections 6.46.7 we assume that $k\geq 2$ — the case $k=1$ (in which the sphere becomes degenerate) requires a separate treatment (although many arguments remain valid for $k=1$) and will be considered in Section 6.8.
6.4. Proof of Theorem 6.1: Conditional expectation kernels
In the present section, we determine the conditional expectation kernels of the $U$-statistic $\mathbb{U}_{n}$. In particular, we shall see that $\zeta_{1}=0$ and $\zeta_{2}>0$.
Recall from Section 6.2 that for $c\in\{1,\ldots,m\}$ and unit vectors $v_{1},\ldots,v_{c}\in\mathbb{S}^{k-1}$, we consider the conditional expectation kernels
$g_{c}(v_{1},\ldots,v_{c})=\mathbb{E}h(v_{1},\ldots,v_{c},V_{c+1},\ldots,V_{m})=\mathbb{P}[\mathop{\mathrm{pos}}\nolimits(v_{1},\ldots,v_{c},V_{c+1},\ldots,V_{m})=\mathbb{R}^{k}].$
Wendel’s formula (1.1) gives
$g_{0}=\mathbb{E}h(V_{1},\ldots,V_{m})=\mathbb{P}[\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{m})=\mathbb{R}^{k}]=p(m,k),\qquad\mathbb{E}\mathbb{U}_{n}=g_{0}=p(m,k).$
The centered version of $g_{c}$ is therefore
$\widetilde{g}_{c}(v_{1},\ldots,v_{c})=g_{c}(v_{1},\ldots,v_{c})-\mathbb{E}h(V_{1},\ldots,V_{m})=g_{c}(v_{1},\ldots,v_{c})-p(m,k).$
Lemma 6.5 ($c=1$).
We have $g_{1}(v_{1})=p(m,k)$ and $\widetilde{g}_{1}(v_{1})=0$ for every $v_{1}\in\mathbb{S}^{k-1}$. Hence $\zeta_{1}=0$.
Proof.
Indeed, by rotation invariance, for every $v_{1}\in\mathbb{S}^{k-1}$ we have
$g_{1}(v_{1})=\mathbb{P}[\mathop{\mathrm{pos}}\nolimits(v_{1},V_{2},\ldots,V_{m})=\mathbb{R}^{k}]=\mathbb{P}[\mathop{\mathrm{pos}}\nolimits(V_{1},V_{2},\ldots,V_{m})=\mathbb{R}^{k}]=p(m,k).$
Hence $\widetilde{g}_{1}(v_{1})=g_{1}(v_{1})-p(m,k)=0$ and $\zeta_{1}=\mathop{\mathrm{Var}}\nolimits\widetilde{g}_{1}(V_{1})=0$. ∎
The next lemma provides an explicit formula for the conditional expectation kernel $g_{c}(v_{1},\ldots,v_{c})$ of arbitrary order. The formula is expressed in terms of conic intrinsic volumes. To each closed convex cone $C\subseteq\mathbb{R}^{k}$ one can associate conic intrinsic volumes $\upsilon_{0}(C),\ldots,\upsilon_{k}(C)$, which are nonnegative and satisfy $\upsilon_{0}(C)+\cdots+\upsilon_{k}(C)=1$. We refer to [1, 2] and [42, Section 2.3] for their definition and properties.
Lemma 6.6 (Conditional expectation kernels).
Let $c\in\{1,\ldots,m-1\}$. Suppose that $v_{1},\ldots,v_{c}\in\mathbb{S}^{k-1}$, $k\geq 2$, are such that the polyhedral cone $C=\mathop{\mathrm{pos}}\nolimits(v_{1},\ldots,v_{c})\subseteq\mathbb{R}^{k}$ is not a linear subspace. Then
(6.11) $g_{c}(v_{1},\ldots,v_{c})=2\sum_{\begin{subarray}{c}i=1,3,5,\ldots\\ i\leq k\end{subarray}}\sum_{j=0}^{k-i}\upsilon_{k-j}(C)\left(\frac{\mathbbm{1}_{\{i+j\neq k\}}}{2^{m-c}}\binom{m-c}{i+j}+\frac{\mathbbm{1}_{\{i+j=k\}}}{2^{m-c}}\sum_{\ell=k}^{m-c}\binom{m-c}{\ell}\right)$,
where $\upsilon_{0}(C),\ldots,\upsilon_{k}(C)$ are the conic intrinsic volumes of the cone $C$.
Proof.
Put $D:=\mathop{\mathrm{pos}}\nolimits(V_{c+1},\ldots,V_{m})\subseteq\mathbb{R}^{k}$. Passing to dual cones, we get
$g_{c}(v_{1},\ldots,v_{c})=\mathbb{P}[\mathop{\mathrm{pos}}\nolimits(C\cup D)=\mathbb{R}^{k}]=\mathbb{P}[C^{\circ}\cap D^{\circ}=\{0\}].$
Let $\mathbb{O}$ be a random $k\times k$-matrix that is uniformly (Haar) distributed on the group $SO(k)$. Additionally, let $\mathbb{O}$ be independent of the $\sigma$-algebra $\mathcal{V}$ generated by $V_{1},\ldots,V_{n}$. Since the law of $D$ (and hence $D^{\circ}$) is invariant under $SO(k)$, the randomly rotated cone $\mathbb{O}D^{\circ}$ has the same distribution as $D^{\circ}$ and we can write
$g_{c}(v_{1},\ldots,v_{c})=\mathbb{P}[C^{\circ}\cap\mathbb{O}D^{\circ}=\{0\}]=\mathbb{E}\Big[\,\mathbb{P}\left[C^{\circ}\cap\mathbb{O}D^{\circ}=\{0\}\,|\,\mathcal{V}\right]\,\Big].$
In the conditional probability, the only random element is the random matrix $\mathbb{O}$. To compute the conditional probability, we invoke the conic kinematic formula whose proof can be found in [1, Corollary 5.2] or in [42, Theorem  4.3.5]; for applications see also [1]. The formula states that for (deterministic) polyhedral cones $A,B\subseteq\mathbb{R}^{k}$ that are not both linear subspaces,
$\mathbb{P}[A\cap\mathbb{O}B=\{0\}]=2\sum_{\begin{subarray}{c}i=1,3,5,\ldots\\ i\leq k\end{subarray}}\upsilon_{k-i}(A\oplus B)=2\sum_{\begin{subarray}{c}i=1,3,5,\ldots\\ i\leq k\end{subarray}}\sum_{j=0}^{k-i}\upsilon_{j}(A)\upsilon_{k-i-j}(B).$
Here, $A\oplus B:=\{(a,b):a\in A,b\in B\}\subseteq\mathbb{R}^{k}\oplus\mathbb{R}^{k}$ denotes the direct (orthogonal) sum of $A$ and $B$. Recall that $v_{1},\ldots,v_{c}\in\mathbb{S}^{k-1}$ are such that $C=\mathop{\mathrm{pos}}\nolimits(v_{1},\ldots,v_{c})$ (and hence $C^{\circ}$) is not a linear subspace. The kinematic formula gives
$\mathbb{P}\left[C^{\circ}\cap\mathbb{O}D^{\circ}=\{0\}\,|\,\mathcal{V}\right]=2\sum_{\begin{subarray}{c}i=1,3,5,\ldots\\ i\leq k\end{subarray}}\sum_{j=0}^{k-i}\upsilon_{j}(C^{\circ})\upsilon_{k-i-j}(D^{\circ})=2\sum_{\begin{subarray}{c}i=1,3,5,\ldots\\ i\leq k\end{subarray}}\sum_{j=0}^{k-i}\upsilon_{k-j}(C)\upsilon_{i+j}(D),$
where we used the duality relation $\upsilon_{j}(C^{\circ})=\upsilon_{k-j}(C)$ and similarly for $D$. Recall that $C$ is deterministic and $D$ is random. Taking expectation gives
(6.12) $g_{c}(v_{1},\ldots,v_{c})=2\sum_{\begin{subarray}{c}i=1,3,5,\ldots\\ i\leq k\end{subarray}}\sum_{j=0}^{k-i}\upsilon_{k-j}(C)\mathbb{E}\upsilon_{i+j}(D).$
Next recall that $D=\mathop{\mathrm{pos}}\nolimits(V_{c+1},\ldots,V_{m})$ has the same law as the random polyhedral cone $\mathcal{W}_{m-c,k}$. To proceed, we need formulas for the expected conic intrinsic volumes of this cone.
Proposition 6.7 (Expected conic intrinsic volumes of $\mathcal{W}_{n,d}$).
For all $d\in\mathbb{N}$, $n\in\mathbb{N}$ we have
(6.13) $\mathbb{E}\left[\upsilon_{k}(\mathcal{W}_{n,d})\right]=\mathbb{E}\left[\upsilon_{k}(\mathcal{W}_{n,d})\mathbbm{1}_{\{\mathcal{W}_{n,d}\neq\mathbb{R}^{d}\}}\right]=\frac{1}{2^{n}}\binom{n}{k},\qquad k\in\{0,\dots,d-1\},$
(6.14) $\mathbb{E}\left[\upsilon_{d}(\mathcal{W}_{n,d})\right]=1-\frac{1}{2^{n}}\sum_{\ell=0}^{d-1}\binom{n}{\ell}=\frac{1}{2^{n}}\sum_{\ell=d}^{n}\binom{n}{\ell}$.
In the special case where $d\in\mathbb{N}$ and $n\in\{1,\ldots,d\}$, these equations state that
(6.15) $\mathbb{E}[\upsilon_{k}(\mathcal{W}_{n,d})]=\frac{1}{2^{n}}\binom{n}{k},\qquad k\in\{0,\dots,n\}.$
Proof.
Equations (6.13) and (6.14) are proved in [15, Lemma 5.1] (building on [20, Corollary 4.3]); where it is assumed that $n\geq d$. See also [24, Theorem 5.18] for an alternative approach. We sketch a proof of (6.15) for $n\in\{1,\dots,d\}$; for background on conic intrinsic volumes and hyperplane arrangements we refer to [42, Section 5.2]. For each sign vector $\varepsilon=(\varepsilon_{1},\dots,\varepsilon_{n})\in\{\pm 1\}^{n}$, consider the random cone $C_{\varepsilon}:=\mathop{\mathrm{pos}}\nolimits(\varepsilon_{1}U_{1},\dots,\varepsilon_{n}U_{n})$. By symmetry, $\mathbb{E}\,\upsilon_{k}(C_{\varepsilon})=\mathbb{E}\,\upsilon_{k}(\mathcal{W}_{n,d})$ for every $\varepsilon$. On the other hand, with probability $1$, the cones $C_{\varepsilon}$ are precisely the chambers of a hyperplane arrangement in $\mathop{\mathrm{lin}}\nolimits(U_{1},\dots,U_{n})$ that has the same intersection poset as the coordinate hyperplane arrangement in $\mathbb{R}^{n}$, whose chambers are the $2^{n}$ orthants. Since the $k$th conic intrinsic volume of an orthant in $\mathbb{R}^{n}$ equals $\frac{1}{2^{n}}\binom{n}{k}$, the Klivans–Swartz formula [42, Theorem 5.1.4] yields $\sum_{\varepsilon\in\{\pm 1\}^{n}}\upsilon_{k}(C_{\varepsilon})=\binom{n}{k}$ a.s. Taking expectations gives $2^{n}\,\mathbb{E}\,\upsilon_{k}(\mathcal{W}_{n,d})=\sum_{\varepsilon\in\{\pm 1\}^{n}}\mathbb{E}\,\upsilon_{k}(C_{\varepsilon})=\binom{n}{k}$, which is precisely (6.15). ∎
Returning to (6.12) and plugging the values of $\mathbb{E}\upsilon_{i+j}(D)=\mathbb{E}\upsilon_{i+j}(\mathcal{W}_{m-c,k})$ given by (6.13) an (6.14) yields
$g_{c}(v_{1},\ldots,v_{c})=2\sum_{\begin{subarray}{c}i=1,3,5,\ldots\\ i\leq k\end{subarray}}\sum_{j=0}^{k-i}\upsilon_{k-j}(C)\left(\frac{\mathbbm{1}_{\{i+j\neq k\}}}{2^{m-c}}\binom{m-c}{i+j}+\frac{\mathbbm{1}_{\{i+j=k\}}}{2^{m-c}}\sum_{\ell=k}^{m-c}\binom{m-c}{\ell}\right)$.
The proof of Lemma 6.6 is complete. ∎
We have already seen that $g_{0}=g_{1}(v_{1})=p(m,k)$ and hence $\widetilde{g}_{1}(v_{1})=0$ for all $v_{1}\in\mathbb{S}^{k-1}$. The next step is to compute $g_{2}(v_{1},v_{2})$.
Lemma 6.8 (Conditional expectation kernel for $c=2$).
Let $v_{1},v_{2}\in\mathbb{S}^{k-1}$, $k\geq 2$, be two vectors such that $v_{1}\neq\pm v_{2}$. Then,
$\widetilde{g}_{2}(v_{1},v_{2})=\frac{1}{2^{m-1}}\binom{m-2}{k-1}\left(\frac{2}{\pi}\arccos\langle v_{1},v_{2}\rangle-1\right)=-\frac{1}{2^{m-1}}\binom{m-2}{k-1}\cdot\frac{2}{\pi}\arcsin\langle v_{1},v_{2}\rangle.$
Proof.
First we compute $g_{2}(v_{1},v_{2})$ by applying Lemma 6.6 with $c=2$. The cone $C=\mathop{\mathrm{pos}}\nolimits(v_{1},v_{2})$ is a wedge with angle $\alpha=\arccos\langle v_{1},v_{2}\rangle$ (measured in radians). Its conic intrinsic volumes are given by $\upsilon_{0}(C)=\frac{1}{2}-\frac{\alpha}{2\pi}$, $\upsilon_{1}(C)=\frac{1}{2}$, $\upsilon_{2}(C)=\frac{\alpha}{2\pi}$, and $\upsilon_{m}(C)=0$ whenever $m\geq 3$. So $\upsilon_{k-j}(C)=0$ unless $k-j\in\{0,1,2\}$. It follows that only the pairs (i,j)=(1,k1)(i,j)=(1,k-1) and (i,j)=(1,k2)(i,j)=(1,k-2) contribute to the sum in (6.11). So
$g_{2}(v_{1},v_{2})$ $=2\upsilon_{1}(C)\cdot\frac{1}{2^{m-2}}\sum_{\ell=k}^{m-2}\binom{m-2}{\ell}+2\upsilon_{2}(C)\cdot\frac{1}{2^{m-2}}\binom{m-2}{k-1}$
$=\frac{1}{2^{m-2}}\sum_{\ell=k}^{m-2}\binom{m-2}{\ell}+\frac{\alpha}{2^{m-2}\pi}\binom{m-2}{k-1}$.
Wendel’s formula (1.1) and the defining property of the Pascal triangle give
$p(m,k)=\mathbb{P}[\mathcal{W}_{m,k}=\mathbb{R}^{k}]=\frac{1}{2^{m-1}}\sum_{\ell=k}^{m-1}\binom{m-1}{\ell}=\frac{1}{2^{m-2}}\sum_{\ell=k}^{m-2}\binom{m-2}{\ell}+\frac{1}{2^{m-1}}\binom{m-2}{k-1}$.
Plugging these formulas into $\widetilde{g}_{2}(v_{1},v_{2})=g_{2}(v_{1},v_{2})-p(m,k)$ gives the stated formula for $\widetilde{g}_{2}(v_{1},v_{2})$. ∎
6.5. Proof of Theorem 6.1: The variance
We have already seen that $\zeta_{1}=\mathop{\mathrm{Var}}\nolimits g_{1}(V_{1})=0$. In the present section, we compute $\zeta_{2}=\mathop{\mathrm{Var}}\nolimits g_{2}(V_{1},V_{2})$. We shall see that $\zeta_{2}>0$, which means that the $U$-statistic $\mathbb{U}_{n}$ is degenerate of order $2$.
Lemma 6.9 (Angle between two random unit vectors).
Let $V_{1},V_{2}$ be independent and uniformly distributed on $\mathbb{S}^{k-1}$, $k\geq 2$. Let $\Theta:=\arccos\langle V_{1},V_{2}\rangle\in[0,\pi]$ be the angle between these vectors. Then,
$\mathbb{E}\Theta=\frac{\pi}{2},\qquad\mathop{\mathrm{Var}}\nolimits\Theta=\frac{1}{2}\psi_{1}\left(\frac{k}{2}\right),$
where $\psi_{1}(x)=\frac{\textup{d}^{2}}{\textup{d}x^{2}}\log\Gamma(x)=\sum_{j=0}^{\infty}\frac{1}{(j+x)^{2}}>0$, with $x>0$, is the trigamma function.
Proof.
It is a standard fact, see, e.g., [26, Example 2.28], that $\Theta$ has density
(6.16) $f_{\Theta}(\theta)=\frac{(\sin\theta)^{k-2}}{Z_{k}},\qquad\theta\in[0,\pi],\qquad Z_{k}=\int_{0}^{\pi}(\sin u)^{k-2}\,\textup{d}u=\frac{\sqrt{\pi}\,\Gamma\!\left(\frac{k-1}{2}\right)}{\Gamma\!\left(\frac{k}{2}\right)},$
for $k\geq 2$. It follows that $\Theta$ has the same law as $\pi-\Theta$, hence $\mathbb{E}\Theta=\pi/2$. Next, it follows that
(6.17) $\mathop{\mathrm{Var}}\nolimits(\Theta)=\mathbb{E}\!\left[\left(\Theta-\frac{\pi}{2}\right)^{2}\right]=\frac{\int_{0}^{\pi}\left(\theta-\frac{\pi}{2}\right)^{2}(\sin\theta)^{k-2}\,\textup{d}\theta}{\int_{0}^{\pi}(\sin\theta)^{k-2}\,\textup{d}\theta}=\frac{\int_{-\pi/2}^{\pi/2}x^{2}\cos^{k-2}x\,\textup{d}x}{\int_{-\pi/2}^{\pi/2}\cos^{k-2}x\,\textup{d}x},$
where we used the substitution $x=\theta-\pi/2$. It remains to show that the quotient on the right-hand side equals $\frac{1}{2}\psi_{1}(k/2)$. We use the identity (see [16, Eq. 3.631.9])
(6.18) $F(a):=\int_{-\pi/2}^{\pi/2}\cos^{k-2}x\,\cos(ax)\,\textup{d}x=\frac{2\,\pi\,\Gamma(k-1)}{2^{k-1}\,\Gamma\!\big(\frac{k}{2}+\tfrac{a}{2}\big)\,\Gamma\!\big(\frac{k}{2}-\tfrac{a}{2}\big)},\qquad k>1,\quad a\in\mathbb{C}.$
On the one hand, the quotient in (6.17) equals $F^{\prime\prime}(0)/F(0)$. On the other hand, taking logarithms in (6.18), differentiating twice, and setting $a=0$ gives $(\log F)^{\prime}(0)=0$ and $(\log F)^{\prime\prime}(0)=-\frac{1}{2}\psi_{1}(k/2)$. Hence, $F^{\prime\prime}(0)/F(0)=-(\log F)^{\prime\prime}(0)=\frac{1}{2}\psi_{1}(k/2)$. ∎
Corollary 6.10 (Formula for $\zeta_{2}$).
For $k\geq 2$ we have
$\zeta_{2}=\frac{2}{2^{2m-2}\pi^{2}}\binom{m-2}{k-1}^{2}\cdot\psi_{1}\left(\frac{k}{2}\right)>0.$
Proof.
Recall that $\zeta_{2}=\mathop{\mathrm{Var}}\nolimits g_{2}(V_{1},V_{2})=\mathop{\mathrm{Var}}\nolimits\widetilde{g}_{2}(V_{1},V_{2})$. Hence
$\zeta_{2}=\mathop{\mathrm{Var}}\nolimits\widetilde{g}_{2}(V_{1},V_{2})=\frac{1}{2^{2m-2}}\binom{m-2}{k-1}^{2}\cdot\frac{4}{\pi^{2}}\mathop{\mathrm{Var}}\nolimits\arccos\langle V_{1},V_{2}\rangle=\frac{2}{2^{2m-2}\pi^{2}}\binom{m-2}{k-1}^{2}\cdot\psi_{1}\left(\frac{k}{2}\right),$
using first Lemma 6.8 and then Lemma 6.9. ∎
We are now able to prove the asymptotic formula for the variance stated in (6.4).
Proposition 6.11 (Asymptotics of the variance).
Let $k\geq 2$ and $q\in\mathbb{N}$ be fixed. Then
$\mathop{\mathrm{Var}}\nolimits\left(\frac{f_{d-q}(\mathcal{W}_{d+k,d})}{\binom{d+k}{d-q}}\right)\sim\frac{\binom{k+q}{2}^{2}\binom{k+q-2}{k-1}^{2}\psi_{1}\left(\frac{k}{2}\right)}{2^{2k+2q-4}\pi^{2}\cdot d^{2}},\qquad d\to\infty.$
Proof.
We have shown that $\zeta_{1}=0$ and $\zeta_{2}>0$. Equation (6.8) with $c=2$ gives
$\mathop{\mathrm{Var}}\nolimits\left(\frac{f_{d-q}(\mathcal{W}_{d+k,d})}{\binom{d+k}{d-q}}\right)=\mathop{\mathrm{Var}}\nolimits\mathbb{U}_{n}\sim\frac{\binom{m}{2}^{2}\,2!}{n^{2}}\;\zeta_{2}=\frac{\binom{m}{2}^{2}\binom{m-2}{k-1}^{2}\psi_{1}\left(\frac{k}{2}\right)}{2^{2m-4}\pi^{2}\cdot n^{2}},\qquad n\to\infty.$
Recalling that $m=k+q$ is fixed and $n=d+k\sim d$ gives the stated asymptotic equivalence. ∎
6.6. Proof of Theorem 6.1: Convolution operator on the sphere and its eigenvalues
Let $k\geq 2$ be integer and recall that $\nu_{k}$ is the uniform probability measure on the unit sphere $\mathbb{S}^{k-1}=\{v\in\mathbb{R}^{k}:\|v\|=1\}$. Define the integral operator $A:L^{2}(\mathbb{S}^{k-1},\nu_{k})\to L^{2}(\mathbb{S}^{k-1},\nu_{k})$ by
(6.19) $(Af)(v)=\int_{\mathbb{S}^{k-1}}\frac{2}{\pi}\arcsin\!\big(\langle v,u\rangle\big)\,f(u)\,\nu_{k}(\textup{d}u),\qquad f\in L^{2}(\mathbb{S}^{k-1},\nu_{k}),\quad v\in\mathbb{S}^{k-1}.$
Note that $A$ is Hilbert–Schmidt (hence compact) since its kernel $K(u,v)=\frac{2}{\pi}\arcsin(\langle v,u\rangle)$ is square integrable w.r.t. $\nu_{k}\otimes\nu_{k}$. Also, $A$ is self-adjoint since $K$ is real-valued and symmetric in its arguments. By the spectral theorem for compact operators, $A$ has discrete spectrum consisting of countably many eigenvalues.
Our aim is to describe the full spectral decomposition of $A$ (i.e. its eigenvalues and the corresponding eigenspaces). This will be done in terms of spherical harmonics. Let us recall some basic facts from the harmonic analysis on the sphere; see [10, Chapters 1,2], [3, Chapter 5], [37]. For $n\in\mathbb{N}_{0}$ let $\mathcal{H}_{n}\subseteq L^{2}(\mathbb{S}^{k-1},\nu_{k})$ be the space of spherical harmonics of degree $n$. One possible definition of $\mathcal{H}_{n}$ is this one:
$\mathcal{H}_{n}=\Big\{Y\in C^{\infty}(\mathbb{S}^{k-1}):\Delta_{\mathbb{S}^{k-1}}Y=-n(n+k-2)\,Y\Big\},$
where $\Delta_{\mathbb{S}^{k-1}}$ denotes the Laplace–Beltrami operator on $\mathbb{S}^{k-1}$. It is known [3, Proposition 5.8 on p. 78] that the linear space $\mathcal{H}_{n}$ is finite-dimensional with
$d_{n,k}:=\dim\mathcal{H}_{n}=\binom{n+k-1}{n}-\binom{n+k-3}{n-2}$.
Moreover, $\mathcal{H}_{m}\perp\mathcal{H}_{n}$ for $m\neq n$, and the linear span of $\bigcup_{n\geq 0}\mathcal{H}_{n}$ is dense in $L^{2}(\mathbb{S}^{k-1},\nu_{k})$; see [10, Theorem 2.2.2]. Note that the excluded case $k=1$ is different since then the Hilbert space $L^{2}(\mathbb{S}^{k-1},\nu_{k})$ becomes $2$-dimensional — see Section 6.8 for this case.
Lemma 6.12 (Spectral decomposition of $A$).
For every $n\in\mathbb{N}_{0}$ and every degree $n$ spherical harmonic $Y\in\mathcal{H}_{n}$, we have $AY=\lambda_{n}\,Y$, where
$\lambda_{n}=0\quad(\text{if $n$ is even}),\qquad\qquad\lambda_{n}=\left(\frac{\Gamma\!\left(\frac{k}{2}\right)\Gamma\!\left(\frac{n}{2}\right)}{\pi\,\Gamma\!\left(\frac{n+k}{2}\right)}\right)^{2}=\frac{\left(B\left(\frac{n}{2},\frac{k}{2}\right)\right)^{2}}{\pi^{2}}\quad(\text{if $n$ is odd}).$
To prove Lemma 6.12, we represent $A$ as the square of another operator.
Lemma 6.13 (Square root of $A$).
We have $A=S^{2}$, where $S:L^{2}(\mathbb{S}^{k-1},\nu_{k})\to L^{2}(\mathbb{S}^{k-1},\nu_{k})$ is a linear operator defined by
$(Sf)(v)=\int_{\mathbb{S}^{k-1}}\operatorname{sign}\!\big(\langle v,u\rangle\big)\,f(u)\,\nu_{k}(\textup{d}u),\qquad f\in L^{2}(\mathbb{S}^{k-1},\nu_{k}),\quad v\in\mathbb{S}^{k-1}.$
Proof.
Let $U$ be uniformly distributed on $\mathbb{S}^{k-1}$. For all $v,w\in\mathbb{S}^{k-1}$ it is easy to check (see, e.g., [49, Lemma 6.7 on p. 144]) that
$\mathbb{E}\!\left[\operatorname{sign}\!\big(\langle v,U\rangle\big)\operatorname{sign}\!\big(\langle U,w\rangle\big)\right]=\frac{2}{\pi}\arcsin(\langle v,w\rangle).$
Applying the definition of $S$ twice gives
$(S^{2}f)(v)=\int_{\mathbb{S}^{k-1}}\!\!\left(\int_{\mathbb{S}^{k-1}}\operatorname{sign}\!\big(\langle v,u\rangle\big)\operatorname{sign}\!\big(\langle u,w\rangle\big)\,\nu_{k}(\textup{d}u)\right)f(w)\,\nu_{k}(\textup{d}w).$
Thus $S^{2}$ is an integral operator with kernel
$L(v,w):=\int_{\mathbb{S}^{k-1}}\operatorname{sign}\!\big(\langle v,u\rangle\big)\operatorname{sign}\!\big(\langle u,w\rangle\big)\,\nu_{k}(\textup{d}u)=\mathbb{E}\!\left[\operatorname{sign}\!\big(\langle v,U\rangle\big)\operatorname{sign}\!\big(\langle U,w\rangle\big)\right]=\frac{2}{\pi}\arcsin(\langle v,w\rangle).$
We conclude that $A=S^{2}$. ∎
The next lemma describes the spectral decomposition of $S$ and, together with the formula $A=S^{2}$, implies Lemma 6.12.
Lemma 6.14 (Spectral decomposition of $S$).
For every $n\in\mathbb{N}_{0}$ and every $Y\in\mathcal{H}_{n}$, we have $SY=\tau_{n}\,Y$ (i.e. the restriction of $S$ to $\mathcal{H}_{n}$ is the multiplication by $\tau_{n}$), where
$\tau_{n}=0\quad(\text{if $n$ is even}),\qquad\qquad\tau_{n}=(-1)^{\frac{n-1}{2}}\,\frac{\Gamma\!\left(\frac{k}{2}\right)\Gamma\!\left(\frac{n}{2}\right)}{\pi\,\Gamma\!\left(\frac{n+k}{2}\right)},\quad(\text{if $n$ is odd}).$
Proof.
The main tool in this proof is the Funk–Hecke formula [10, Theorem 1.2.9] which we now recall. Let $\alpha=\frac{k-2}{2}\geq 0$ and let $h:[-1,1]\to\mathbb{C}$ be integrable with respect to the measure $(1-t^{2})^{\alpha-\frac{1}{2}}\,\textup{d}t$. Define the integral operator $T_{h}:L^{2}(\mathbb{S}^{k-1},\nu_{k})\to L^{2}(\mathbb{S}^{k-1},\nu_{k})$ by
$(T_{h}f)(v):=\int_{\mathbb{S}^{k-1}}h(\langle v,u\rangle)\,f(u)\,\nu_{k}(\textup{d}u).$
The Funk–Hecke formula states that for every $n\in\mathbb{N}_{0}$ and every degree $n$ spherical harmonic $Y\in\mathcal{H}_{n}$, we have $T_{h}Y=\lambda_{n}(h)\,Y$, where
$\lambda_{n}(h)=\frac{b_{\alpha}}{C_{n}^{(\alpha)}(1)}\int_{-1}^{1}h(t)\,C_{n}^{(\alpha)}(t)\,(1-t^{2})^{\alpha-\frac{1}{2}}\,\textup{d}t,\qquad b_{\alpha}=\left(\int_{-1}^{1}(1-t^{2})^{\alpha-\frac{1}{2}}\,\textup{d}t\right)^{-1}=\frac{\Gamma(\alpha+1)}{\sqrt{\pi}\,\Gamma(\alpha+\tfrac{1}{2})}.$
Here $C_{n}^{(\alpha)}$ denotes the Gegenbauer polynomial [10, Sections B.1,B.2] of degree $n$. For our purposes, it is convenient to define it by the Rodrigues’ formula (see [10, Equations (B.1.2), (B.2.1)])
$C_{n}^{(\alpha)}(t)\,(1-t^{2})^{\alpha-\frac{1}{2}}=\kappa_{n}^{(\alpha)}\,\frac{\textup{d}^{n}}{\textup{d}t^{n}}\Big((1-t^{2})^{n+\alpha-\frac{1}{2}}\Big),\qquad\kappa_{n}^{(\alpha)}=\frac{(-1)^{n}}{2^{n}n!}\,\frac{\Gamma\!\left(\alpha+\tfrac{1}{2}\right)\Gamma(n+2\alpha)}{\Gamma(2\alpha)\Gamma\!\left(n+\alpha+\tfrac{1}{2}\right)}.$
In our setting, $h(t)=\operatorname{sign}(t)$, and the Funk–Hecke formula gives $SY=\tau_{n}\,Y$ for all $Y\in\mathcal{H}_{n}$, where
$\tau_{n}=\frac{b_{\alpha}}{C_{n}^{(\alpha)}(1)}\,I_{n}^{(\alpha)}\qquad\text{ with }\qquad I_{n}^{(\alpha)}:=\int_{-1}^{1}\operatorname{sign}(t)\,C_{n}^{(\alpha)}(t)\,(1-t^{2})^{\alpha-\frac{1}{2}}\,\textup{d}t.$
It remains to compute $I_{n}^{(\alpha)}$. Writing $F_{n}(t):=(1-t^{2})^{n+\alpha-\frac{1}{2}}$, splitting the integral and using the fact that the antiderivative of $F_{n}^{(n)}$ is $F_{n}^{(n-1)}$ gives
$I_{n}^{(\alpha)}=\kappa_{n}^{(\alpha)}\int_{-1}^{1}\operatorname{sign}(t)\,F_{n}^{(n)}(t)\,\textup{d}t=-\kappa_{n}^{(\alpha)}F_{n}^{(n-1)}(t)\big|_{t=-1}^{t=0}+\kappa_{n}^{(\alpha)}F_{n}^{(n-1)}(t)\big|_{t=0}^{t=1}.$
Since the vanishing order of $F(t)$ at $t=\pm 1$ is $n+\alpha-\frac{1}{2}>n-1$, we have $F_{n}^{(n-1)}(\pm 1)=0$. It follows that
$I_{n}^{(\alpha)}=-2\kappa_{n}^{(\alpha)}\,F_{n}^{(n-1)}(0).$
Taking these results together gives
(6.20) $\tau_{n}=\frac{b_{\alpha}}{C_{n}^{(\alpha)}(1)}\,I_{n}^{(\alpha)}=\frac{b_{\alpha}}{C_{n}^{(\alpha)}(1)}\,\cdot(-2\kappa_{n}^{(\alpha)})F_{n}^{(n-1)}(0)=\frac{(-1)^{n+1}}{2^{n-1}}\,\frac{\Gamma(\alpha+1)}{\sqrt{\pi}}\,\frac{F_{n}^{(n-1)}(0)}{\Gamma(n+\alpha+\frac{1}{2})},$
where we used that
$\kappa_{n}^{(\alpha)}=\frac{(-1)^{n}}{2^{n}n!}\,\frac{\Gamma\!\left(\alpha+\frac{1}{2}\right)\Gamma(n+2\alpha)}{\Gamma(2\alpha)\Gamma\!\left(n+\alpha+\frac{1}{2}\right)},\qquad b_{\alpha}=\frac{\Gamma(\alpha+1)}{\sqrt{\pi}\,\Gamma\!\left(\alpha+\frac{1}{2}\right)},\qquad C_{n}^{(\alpha)}(1)=\frac{\Gamma(n+2\alpha)}{\Gamma(2\alpha)\,n!};$
see [10, Equation (B.2.2)] for the formula for $C_{n}^{(\alpha)}(1)$. The function $F(t)$ is even and (6.20) implies that $\tau_{n}=0$ for even $n$. Let $n$ be odd. Expanding around $t=0$ gives
$F_{n}(t)=(1-t^{2})^{\,n+\alpha-\frac{1}{2}}=\sum_{j=0}^{\infty}(-1)^{j}\binom{n+\alpha-\frac{1}{2}}{j}\,t^{2j}.$
Then n1n-1 is even and the term contributing to $t^{n-1}$ is $j=\frac{n-1}{2}$, hence
$F_{n}^{(n-1)}(0)=(n-1)!\,(-1)^{\frac{n-1}{2}}\binom{n+\alpha-\frac{1}{2}}{\frac{n-1}{2}}=(-1)^{\frac{n-1}{2}}\,\frac{\Gamma(n)\,\Gamma\!\left(n+\alpha+\frac{1}{2}\right)}{\Gamma\!\left(\frac{n+1}{2}\right)\,\Gamma\!\left(\alpha+\frac{n+2}{2}\right)}.$
Plugging this into (6.20) and using Legendre’s duplication formula $\Gamma(n)=2^{\,n-1}\,\frac{\Gamma\!\left(\frac{n}{2}\right)\,\Gamma\!\left(\frac{n+1}{2}\right)}{\sqrt{\pi}}$ gives
$\tau_{n}=(-1)^{\frac{n-1}{2}}\,\frac{\Gamma(\alpha+1)\,\Gamma\!\left(\frac{n}{2}\right)}{\pi\,\Gamma\!\left(\alpha+\frac{n}{2}+1\right)}=(-1)^{\frac{n-1}{2}}\,\frac{\Gamma\!\left(\frac{k}{2}\right)\Gamma\!\left(\frac{n}{2}\right)}{\pi\,\Gamma\!\left(\frac{n+k}{2}\right)},\qquad n\ \text{odd}.$
In the last step, we used $\alpha=\frac{k-2}{2}$. ∎
Proposition 6.15 (Traces of $A$ and $A^{2}$).
The operator $A$ defined in (6.19) is of trace class and
$\operatorname{tr}(A)$ $=\sum_{n=1,3,5,\ldots}d_{n,k}\,\left(\frac{\Gamma\!\left(\frac{k}{2}\right)\Gamma\!\left(\frac{n}{2}\right)}{\pi\,\Gamma\!\left(\frac{n+k}{2}\right)}\right)^{2}\;=\;1,$
$\operatorname{tr}(A^{2})$ $=\sum_{n=1,3,5,\ldots}d_{n,k}\,\left(\frac{\Gamma\!\left(\frac{k}{2}\right)\Gamma\!\left(\frac{n}{2}\right)}{\pi\,\Gamma\!\left(\frac{n+k}{2}\right)}\right)^{4}=\frac{2}{\pi^{2}}\,\psi_{1}\!\left(\frac{k}{2}\right),$
where $d_{n,k}=\dim\mathcal{H}_{n}=\binom{n+k-1}{n}-\binom{n+k-3}{n-2}$ and $\psi_{1}$ is the trigamma function.
Proof.
Let us first show that $A,A^{2}$ are trace class operators and $\operatorname{tr}(A)=1$, $\operatorname{tr}(A^{2})=\frac{2}{\pi^{2}}\psi_{1}(\frac{k}{2})$. Recall that $A=S^{2}=SS^{*}$ and $S=S^{*}$ is Hilbert–Schmidt. It follows that $A$ is trace class and $\operatorname{tr}(A)=\|S\|^{2}_{\text{HS}}$ is the squared Hilbert–Schmidt norm of $S$. Since the kernel of $S$ takes only values $\pm 1$, we have $\|S\|_{\text{HS}}^{2}=1$ and hence $\operatorname{tr}(A)=1$. Similarly, $A^{2}=AA^{*}$ and $A=A^{*}$ is Hilbert–Schmidt, hence $A^{2}$ is of trace class and
$\operatorname{tr}(A^{2})$ $=\|A\|_{\text{HS}}^{2}=\frac{4}{\pi^{2}}\iint_{\mathbb{S}^{k-1}\times\mathbb{S}^{k-1}}\arcsin^{2}\!\big(\langle v,u\rangle\big)\,\nu_{k}(\textup{d}v)\,\nu_{k}(\textup{d}u)=\frac{4}{\pi^{2}}\int_{\mathbb{S}^{k-1}}\arcsin^{2}\!\big(\langle v,e_{1}\rangle\big)\,\nu_{k}(\textup{d}v)$
$=\frac{4}{\pi^{2}}\,\frac{\Gamma\!\left(\frac{k}{2}\right)}{\sqrt{\pi}\,\Gamma\!\left(\frac{k-1}{2}\right)}\int_{-1}^{1}\arcsin^{2}(t)\,(1-t^{2})^{\frac{k-3}{2}}\,\textup{d}t=\frac{4}{\pi^{2}}\,\frac{\Gamma\!\left(\frac{k}{2}\right)}{\sqrt{\pi}\,\Gamma\!\left(\frac{k-1}{2}\right)}\int_{-\pi/2}^{\pi/2}x^{2}\,\cos^{k-2}x\,\textup{d}x,$
where the last step follows by the substitution $t=\sin x$. We have already computed the right-hand side in the proof of Lemma 6.9: it equals $\frac{4}{\pi^{2}}\mathop{\mathrm{Var}}\nolimits\Theta=\frac{2}{\pi^{2}}\psi_{1}(\frac{k}{2})$. So, $\operatorname{tr}(A^{2})=\frac{2}{\pi^{2}}\psi_{1}(\frac{k}{2})$.
On the other hand, we have seen in Lemma 6.12 that the nonzero eigenvalues of $A$ are
$\lambda_{n}=\left(\frac{\Gamma\!\left(\frac{k}{2}\right)\Gamma\!\left(\frac{n}{2}\right)}{\pi\,\Gamma\!\left(\frac{n+k}{2}\right)}\right)^{2}\qquad(\text{with multiplicity }d_{n,k}=\dim\mathcal{H}_{n}),\quad n=1,3,5,\ldots.$
Note that 0 is also an eigenvalue of $A$, and the eigenspace of 0 is the closure of $\oplus_{n=0,2,4,\ldots}\mathcal{H}_{n}$, however, 0 does not contribute to the trace. It follows that $\operatorname{tr}(A)=\sum_{n=1,3,5,\ldots}\lambda_{n}d_{n,k}$ and $\operatorname{tr}(A^{2})=\sum_{n=1,3,5,\ldots}\lambda_{n}^{2}d_{n,k}$, which completes the proof. ∎
Example 6.16 ($k=2$).
For $k=2$ the underlying space is the unit circle $\mathbb{S}^{1}$. We represent $v\in\mathbb{S}^{1}$ as $v=(\cos\theta,\sin\theta)$. The spaces of spherical harmonics are as follows: $\mathcal{H}_{0}=\mathop{\mathrm{lin}}\nolimits\{1\}$ and $\mathcal{H}_{n}=\mathop{\mathrm{lin}}\nolimits\{\cos(n\theta),\sin(n\theta)\}$ for $n\geq 1$. So $\dim\mathcal{H}_{0}=1$ and $\dim\mathcal{H}_{n}=2$ for $n\geq 1$. The non-zero eigenvalues of the operator $A$ are $4/(\pi^{2}n^{2})$, $n=1,3,5,\dots$, each with multiplicity $2$. This can be verified by direct computation.
6.7. Proof of Theorem 6.1: Completing the argument
We now gathered all ingredients needed to prove Theorem 6.1 for $k\geq 2$. Recall from (6.10) the representation
$f_{d-q}(\mathcal{W}_{d+k,d})/\binom{d+k}{d-q}=\mathbb{U}_{n},$
where $\mathbb{U}_{n}$ is the $U$-statistic corresponding to the kernel $h(v_{1},\ldots,v_{m})=\mathbbm{1}_{\{\mathop{\mathrm{pos}}\nolimits(v_{1},\ldots,v_{m})=\mathbb{R}^{k}\}}$, for all $v_{1},\ldots,v_{m}\in\mathbb{S}^{k-1}$. In Lemma 6.5 we have shown that $\zeta_{1}=0$ and in Lemma 6.8 we computed
$\widetilde{g}_{2}(v_{1},v_{2})=-\frac{1}{2^{m-1}}\binom{m-2}{k-1}\cdot\frac{2}{\pi}\arcsin\langle v_{1},v_{2}\rangle,\qquad v_{1}\neq\pm v_{2}.$
In Corollary 6.10 we have seen that $\zeta_{2}>0$. We are thus in the setting of Theorem 6.4. The operator $T:L^{2}(\mathbb{S}^{k-1},\nu_{k})\to L^{2}(\mathbb{S}^{k-1},\nu_{k})$ appearing in Theorem 6.4 is
$(Tf)(x):=\int_{\mathbb{S}^{k-1}}\widetilde{g}_{2}(x,y)\,f(y)\,\nu_{k}(\textup{d}y)=-\frac{1}{2^{m-1}}\binom{m-2}{k-1}\cdot(Af)(x),$
where $A$ is the operator we analyzed in Section 6.6. We have seen in Lemma 6.12 that the non-zero eigenvalues of the operator $A$ are
$\lambda_{r}=\left(\frac{\Gamma\!\left(\frac{k}{2}\right)\Gamma\!\left(\frac{r}{2}\right)}{\pi\,\Gamma\!\left(\frac{r+k}{2}\right)}\right)^{2}\text{ with multiplicity }d_{r,k}=\binom{r+k-1}{r}-\binom{r+k-3}{r-2},\qquad r=1,3,5,\ldots.$
The non-zero eigenvalues of $T$ are then $\frac{1}{2^{m-1}}\binom{m-2}{k-1}\,\lambda_{r}$ with $r=1,3,5,\ldots$ and the same multiplicities $d_{r,k}$ as before. Now we have everything to apply Theorem 6.4. Theorem 6.4 gives
$n\,(\mathbb{U}_{n}-\mathbb{E}\mathbb{U}_{n})\overset{w}{\underset{n\to\infty}{\longrightarrow}}-\frac{\binom{m}{2}\binom{m-2}{k-1}}{2^{m-1}}\,\sum_{r=1,3,5,\ldots}\sum_{j=1}^{d_{r,k}}\lambda_{r}\,(\xi_{r,j}^{2}-1),$
where $\xi_{r,j}$ are i.i.d. standard normal random variables indexed by $r\in\{1,3,5,\ldots\}$ and $j\in\{1,\ldots,d_{r,k}\}$. In Lemma 6.15 we have seen that $\operatorname{tr}(A)=\sum_{r=1,3,5,\ldots}^{\infty}d_{r,k}\lambda_{r}=1$, which allows us to rewrite the above result as
[…]
Now observe that $\sum_{j=1}^{d_{r,k}}\xi_{r,j}^{2}$ is distributed as $\operatorname{Gamma}(\frac{d_{r,k}}{2},\frac{1}{2})$ (chi squared distribution) and these variables are independent for different $r$’s. This proves the distributional representation for $Q_{k}$ stated in Theorem 6.1. By Lemma 6.15, $\mathbb{E}Q_{k}=\operatorname{tr}(A)=1$. Similarly, the variance of $Q_{k}$ is given by
$\mathop{\mathrm{Var}}\nolimits Q_{k}=\sum_{r=1,3,5,\ldots}\lambda_{r}^{2}\,\sum_{j=1}^{d_{r,k}}\mathop{\mathrm{Var}}\nolimits(\xi_{r,j}^{2})=\operatorname{tr}(A^{2})\mathop{\mathrm{Var}}\nolimits(\xi_{1,1}^{2})=\frac{2}{\pi^{2}}\,\psi_{1}\!\left(\frac{k}{2}\right)\mathop{\mathrm{Var}}\nolimits(\xi_{1}^{2})=\frac{4}{\pi^{2}}\,\psi_{1}\!\left(\frac{k}{2}\right),$
where the formula for $\operatorname{tr}(A^{2})$ comes from Lemma 6.15. The asymptotic formula for $\mathop{\mathrm{Var}}\nolimits\mathbb{U}_{n}$ stated in (6.4) has been already established in Proposition 6.11. The proof of Theorem 6.1 in the case where $k\geq 2$ is complete.
6.8. Proof of Theorem 6.1: Case $k=1$
The case where $k=1$ requires special treatment since the unit sphere becomes degenerate, $\mathbb{S}^{0}=\{\pm 1\}$, and the parts of the proof based on the conic kinematic formula and spherical harmonics do not apply directly. Our setting is as follows:
  • The sample space is $\mathcal{X}=\mathbb{S}^{0}=\{\pm 1\}$, endowed with the uniform distribution $\nu_{1}$.
  • $V_{1},V_{2},\ldots,V_{n}$, where $n=d+1$, are i.i.d. with $\mathbb{P}[V_{i}=1]=\mathbb{P}[V_{i}=-1]=1/2$.
  • The order of the $U$-statistic is $m:=n-\ell=q+1$ and the kernel $h:\{\pm 1\}^{m}\to\{0,1\}$ is
    $h(v_{1},\ldots,v_{m}):=\mathbbm{1}_{\{\mathop{\mathrm{pos}}\nolimits(v_{1},\ldots,v_{m})=\mathbb{R}\}}=1-\mathbbm{1}_{\{v_{1}=\ldots=v_{m}\}}.$
The associated $U$-statistic is $\mathbb{U}_{n}=f_{d-q}(\mathcal{W}_{d+1,d})/\binom{d+1}{d-q}$; see (6.10).
Lemma 6.17 (Conditional expectation kernels for $k=1$).
We have $g_{0}=\mathbb{E}\mathbb{U}_{n}=1-2^{1-m}$, $\widetilde{g}_{1}(x)=0$ for $x\in\{\pm 1\}$ and
(6.21) $\widetilde{g}_{2}(x,y)=\begin{cases}+2^{1-m},&x\neq y,\\ -2^{1-m},&x=y,\end{cases}\qquad x,y\in\{\pm 1\}.$
Consequently, $\zeta_{1}=0$ and $\zeta_{2}=2^{2-2m}$.
Proof.
Clearly,
$g_{0}=\mathbb{E}\mathbb{U}_{n}=\mathbb{E}h(V_{1},\ldots,V_{m})=1-\mathbb{P}[V_{1}=\ldots=V_{m}]=1-2\cdot 2^{-m}.$
Next, for $x\in\{\pm 1\}$,
$g_{1}(x)=\mathbb{E}\big[h(x,V_{2},\ldots,V_{m})\big]=1-\mathbb{P}\big[x,V_{2},\ldots,V_{m}\text{ are all equal}\big]=1-2^{-(m-1)}=g_{0}.$
Hence $\widetilde{g}_{1}(x)=g_{1}(x)-g_{0}=0$ and $\zeta_{1}=0$. Next, for $x,y\in\{\pm 1\}$,
$g_{2}(x,y)=\mathbb{E}\big[h(x,y,V_{3},\ldots,V_{m})\big]=1-\mathbb{P}\big[x,y,V_{3},\ldots,V_{m}\text{ are all equal}\big].$
If $x\neq y$, then the event “all equal” is impossible and hence $g_{2}(x,y)=1$. If $x=y$, then “all equal” occurs iff $V_{3}=\cdots=V_{m}=x$, which has probability $2^{-(m-2)}$. Therefore
$g_{2}(x,y)=\begin{cases}1,&x\neq y,\\[2.84526pt] 1-2^{-(m-2)},&x=y,\end{cases}\qquad x,y\in\{\pm 1\}.$
Hence the centered kernel $\widetilde{g}_{2}(x,y)=g_{2}(x,y)-g_{0}=g_{2}(x,y)-1+2\cdot 2^{-m}$ is given by (6.21). Finally, $\zeta_{2}=\mathbb{E}[\widetilde{g}_{2}(V_{1},V_{2})^{2}]=(2^{1-m})^{2}=2^{2-2m}.$
As a linear space, $L^{2}(\{\pm 1\},\nu_{1})$ can be identified with $\mathbb{R}^{2}$ by identifying a function $f$ with the vector $(f(+1),f(-1))^{\top}$. The integral operator $T:L^{2}(\{\pm 1\},\nu_{1})\to L^{2}(\{\pm 1\},\nu_{1})$ with kernel $\widetilde{g}_{2}(x,y)$ appearing in Theorem 6.4 is then represented by the matrix
$T\;\equiv\;2^{-m}\begin{pmatrix}-1&\phantom{-}1\\ \phantom{-}1&-1\end{pmatrix}$.
In particular, $T$ has eigenvalues 0 and $2^{1-m}$. Recall that $n=d+1\sim d$ and $m=q+1$. Theorem 6.4 yields
$n\cdot\left(\frac{f_{d-q}(\mathcal{W}_{d+1,d})-\mathbb{E}f_{d-q}(\mathcal{W}_{d+1,d})}{\binom{d+1}{d-q}}\right)=n\,(\mathbb{U}_{n}-\mathbb{E}\mathbb{U}_{n})\$$\overset{w}{\underset{n\to\infty}{\longrightarrow}}\ -\binom{m}{2}\,2^{1-m}\,(\xi^{2}-1)$
$=\binom{q+1}{2}\,2^{-q}\,(1-\xi^{2}),$
where $\xi$ is standard normal and hence $Q_{1}:=\xi^{2}\stackrel{{\scriptstyle d}}{{=}}\operatorname{Gamma}(\frac{1}{2},\frac{1}{2})$. Applying (6.8) and recalling that $\zeta_{2}=2^{-2q}$ gives
$\mathop{\mathrm{Var}}\nolimits\left(\frac{f_{d-q}(\mathcal{W}_{d+1,d})}{\binom{d+1}{d-q}}\right)=\mathop{\mathrm{Var}}\nolimits\mathbb{U}_{n}\sim\frac{2\,\binom{m}{2}^{2}}{n^{2}}\;\zeta_{2}\sim\frac{\binom{q+1}{2}^{2}}{d^{2}}\;2^{1-2q}.$
Note finally that $\mathbb{E}[\xi^{2}]=1$ and $\mathop{\mathrm{Var}}\nolimits[\xi^{2}]=2=\frac{4}{\pi^{2}}\,\psi_{1}(\frac{1}{2})$. This proves Theorem 6.1 for $k=1$.
6.9. Exponential concentration of face counts
Let us finally mention the following non-asymptotic concentration result. Recall that $p(n,k)$ is given by Wendel’s formula (1.1).
Theorem 6.18 (Exponential concentration of face counts).
Let $d\in\mathbb{N}$, $n>d$, and $\ell\in\{0,\ldots,d-1\}$. Then, for all $t>0$,
$\mathbb{P}\left[\left|\frac{f_{\ell}(\mathcal{W}_{n,d})}{\binom{n}{\ell}}-p(n-\ell,n-d)\right|\geq t\right]\leq 2{\rm e}^{-2t^{2}\lfloor n/(n-\ell)\rfloor}.$
For the proof we need a result of Hoeffding [19] which can be found in the book of Serfling [45, Theorem A on p. 201] and in [39].
Lemma 6.19 (Hoeffding bound for $U$-statistics).
Let the kernel $h:\mathcal{X}^{m}\to\mathbb{R}$ be bounded, that is
$L:=\sup_{x\in\mathcal{X}^{m}}h(x)-\inf_{x\in\mathcal{X}^{m}}h(x)<\infty.$
Then for every $n\geq m$ and $t>0$, we have $\mathbb{P}\left[\left|\mathbb{U}_{n}-\mathbb{E}\mathbb{U}_{n}\right|\geq t\right]\;\leq\;2{\rm e}^{-2\lfloor n/m\rfloor\,t^{2}/L^{2}}.$
Proof of Theorem 6.18.
Recall from Section 6.3 (in particular, Equation (6.9)) the identification
$\mathbb{U}_{n}=\binom{n}{n-\ell}^{-1}\sum_{1\leq i_{1}<\cdots<i_{n-\ell}\leq n}\mathbbm{1}_{\{\mathop{\mathrm{pos}}\nolimits(V_{i_{1}},\ldots,V_{i_{n-\ell}})=\mathbb{R}^{n-d}\}}=\frac{f_{\ell}(\mathcal{W}_{n,d})}{\binom{n}{\ell}}$,
where $X_{i}:=V_{i}$, $i\in\mathbb{N}$, are i.i.d. random elements with the uniform distribution on $\mathcal{X}:=\mathbb{S}^{n-d-1}$. By Proposition 3.2, we have $\mathbb{E}\mathbb{U}_{n}=p(n-\ell,n-d)$. Lemma 6.19 gives
$\mathbb{P}\left[\left|\frac{f_{\ell}(\mathcal{W}_{n,d})}{\binom{n}{\ell}}-p(n-\ell,n-d)\right|\geq t\right]=\mathbb{P}\left[\left|\mathbb{U}_{n}-\mathbb{E}\mathbb{U}_{n}\right|\geq t\right]\;\leq\;2{\rm e}^{-2\lfloor n/m\rfloor\,t^{2}/L^{2}}=2{\rm e}^{-2t^{2}\lfloor n/(n-\ell)\rfloor},$
since $h$ takes values $0,1$ only and hence $L=\sup_{x\in\mathcal{X}^{m}}h(x)-\inf_{x\in\mathcal{X}^{m}}h(x)=1$. ∎
Appendix A Basic facts on Gale duality
In this appendix we collect a number of standard facts about (linear) Gale duality. Although these results are well known, we include proofs for completeness.
We first fix notation and briefly recall linear Gale duality. Let $d^{\prime},d^{\prime\prime}\in\mathbb{N}$ and set $n:=d^{\prime}+d^{\prime\prime}$. Consider vector configurations
$a_{1},\dots,a_{n}\in\mathbb{R}^{d^{\prime}}\qquad\text{and}\qquad b_{1},\dots,b_{n}\in\mathbb{R}^{d^{\prime\prime}}.$
Let $A\in\mathbb{R}^{d^{\prime}\times n}$ be the matrix with columns $a_{1},\dots,a_{n}$ and $B\in\mathbb{R}^{d^{\prime\prime}\times n}$ the matrix with columns $b_{1},\dots,b_{n}$. The row space of $A$, denoted by $\operatorname{Row}(A)$, is the linear subspace of $\mathbb{R}^{n}$ spanned by the rows of $A$; the row space $\operatorname{Row}(B)$ is defined analogously. We recall from Section 2.1 that the configurations $a_{1},\dots,a_{n}$ and $b_{1},\dots,b_{n}$ are in linear Gale duality if
(A.1) $\dim\operatorname{Row}(A)=d^{\prime},\qquad\dim\operatorname{Row}(B)=d^{\prime\prime},\qquad\operatorname{Row}(A)=\operatorname{Row}(B)^{\perp}.$
We begin with a basic criterion for when a finitely generated cone is in fact a linear subspace, phrased in terms of a strictly positive linear dependence among its generators.
Lemma A.1 (Criterion for $\mathop{\mathrm{pos}}\nolimits=\mathop{\mathrm{lin}}\nolimits$).
Let $v_{1},\dots,v_{m}\in\mathbb{R}^{n}$ be vectors. Then the following are equivalent:
  1. (i)
    $\mathop{\mathrm{pos}}\nolimits(v_{1},\dots,v_{m})=\mathop{\mathrm{lin}}\nolimits(v_{1},\dots,v_{m})$.
  2. (ii)
    There exist scalars $\lambda_{1}>0,\dots,\lambda_{m}>0$ such that $\lambda_{1}v_{1}+\cdots+\lambda_{m}v_{m}=0.$
Proof.
[…]
$v_{k}=\sum_{j=1}^{m}\mu_{j;k}\,v_{j}.$
Summing these identities over $k=1,\dots,m$ and rearranging yields $\sum_{j=1}^{m}\lambda_{j}v_{j}=0$ with $\lambda_{j}:=1+\sum_{k=1}^{m}\mu_{j;k}\geq 1$. Hence $\lambda_{1}>0,\dots,\lambda_{m}>0$, proving (ii).
[…]
$v_{k}=\sum_{j:j\neq k}(\lambda_{j}/\lambda_{k})\,v_{j}\in\mathop{\mathrm{pos}}\nolimits(v_{1},\dots,v_{m}),$
implying $\pm v_{k}\in\mathop{\mathrm{pos}}\nolimits(v_{1},\dots,v_{m})$ for every $k$. Therefore $\mathop{\mathrm{lin}}\nolimits(v_{1},\dots,v_{m})\subseteq\mathop{\mathrm{pos}}\nolimits(v_{1},\dots,v_{m})$. The reverse inclusion $\mathop{\mathrm{pos}}\nolimits(v_{1},\dots,v_{m})\subseteq\mathop{\mathrm{lin}}\nolimits(v_{1},\dots,v_{m})$ always holds, proving (i). ∎
The next lemma makes precise how faces of the cone $\mathop{\mathrm{pos}}\nolimits(a_{1},\dots,a_{n})$ correspond to strictly positive dependent subsets (those for which $\mathop{\mathrm{pos}}\nolimits=\mathop{\mathrm{lin}}\nolimits$) in a Gale-dual configuration. It can be found in [46, Theorem 1 on p. 167] or [34, 2A11 on p. 95].
Lemma A.2 (Faces vs. $\mathop{\mathrm{pos}}\nolimits=\mathop{\mathrm{lin}}\nolimits$).
Let $d^{\prime},d^{\prime\prime}\in\mathbb{N}$ and put $n:=d^{\prime}+d^{\prime\prime}$. Let $a_{1},\dots,a_{n}\in\mathbb{R}^{d^{\prime}}$ and $b_{1},\dots,b_{n}\in\mathbb{R}^{d^{\prime\prime}}$ be in linear Gale duality, and set $C_{A}:=\mathop{\mathrm{pos}}\nolimits(a_{1},\dots,a_{n})$. For $I\subseteq[n]$, the following are equivalent:
  1. (i)
    There exists a face $F$ of $C_{A}$ such that $I=\{i\in[n]:a_{i}\in F\}$.
  2. (ii)
    $\mathop{\mathrm{pos}}\nolimits(b_{j}:j\in I^{c})=\mathop{\mathrm{lin}}\nolimits(b_{j}:j\in I^{c})$. Here $\II^{c}:=[n]\backslash I$ is the complement of $I$.
Proof.
Let $A\in\mathbb{R}^{d^{\prime}\times n}$ and $B\in\mathbb{R}^{d^{\prime\prime}\times n}$ be the matrices with columns $a_{1},\dots,a_{n}$ and $b_{1},\dots,b_{n}$, respectively.
Step 1. Every face of $C_{A}$ is of the form $F=\{y\in C_{A}:\langle u,y\rangle=0\}$ for some $u\in\mathbb{R}^{d^{\prime}}$ such that $\langle u,y\rangle\geq 0$ for all $y\in C_{A}$. (When $F=C_{A}$, we may take $u=0$.) Thus, (i) holds if and only if there exists $u\in\mathbb{R}^{d^{\prime}}$ such that
$\langle u,a_{i}\rangle=0\quad(\text{for }i\in I)\qquad\text{and}\qquad\langle u,a_{j}\rangle>0\quad(\text{for }j\in I^{c}).$
With $x:=u^{\top}A\in\mathbb{R}^{n}$, (i) is equivalent to the existence of a vector
(A.2) $x\in\operatorname{Row}(A)\quad\text{ with }\quad x_{i}=0\quad(\text{for }i\in I),\quad x_{j}>0\quad(\text{for }j\in I^{c}).$
Step 2. Apply Lemma A.1 to the vectors $b_{j},j\in I^{c}$. Then (ii) is equivalent to the existence of coefficients $x_{j}>0$ for $j\in I^{c}$ such that $\sum_{j\in I^{c}}x_{j}b_{j}=0$. Extending by $x_{i}:=0$ for $i\in I$ and defining $x=(x_{1},\ldots,x_{n})$, we see that (ii) is equivalent to the existence of a vector
(A.3) $x\in\operatorname{Row}(B)^{\perp}\quad\text{ with }\quad x_{i}=0\quad(\text{for }i\in I),\quad x_{j}>0\quad(\text{for }j\in I^{c}).$
Step 3. By (A.1), $\operatorname{Row}(A)=\operatorname{Row}(B)^{\perp}$, proving the equivalence of (A.2) and (A.3). Hence (i) and (ii) are equivalent. ∎
Recall that pairwise distinct vectors $v_{1},\dots,v_{m}\in\mathbb{R}^{d}$, where $m\geq d$, are in general linear position if every $d$-element subset of $\{v_{1},\dots,v_{m}\}$ is linearly independent.
Lemma A.3.
Let $d^{\prime},d^{\prime\prime}\in\mathbb{N}$ and put $n:=d^{\prime}+d^{\prime\prime}$. Let $a_{1},\dots,a_{n}\in\mathbb{R}^{d^{\prime}}$ and $b_{1},\dots,b_{n}\in\mathbb{R}^{d^{\prime\prime}}$ be in linear Gale duality. If one of the configurations is in general linear position, then the other configuration is in general linear position as well.
Proof.
Assume that $a_{1},\dots,a_{n}$ are in general linear position. Let $J\subseteq[n]$ with $J=d^{\prime\prime}$ and put $I:=J^{c}$, so $I=d^{\prime}$. To show that $b_{j},j\in J$, are linearly independent, let $x\in\mathbb{R}^{n}$ satisfy $x_{1}b_{1}+\ldots+x_{n}b_{n}=0$ and $x_{i}=0$ for all $i\in I$ (so $x$ is supported on $J$). Thus $x\in\operatorname{Row}(B)^{\perp}$ and, by (A.1), $x\in\operatorname{Row}(A)$. Let $A_{I}$ be the $(d^{\prime}\times d^{\prime})$-submatrix of $A$ consisting of the columns indexed by $I$. Since $I=d^{\prime}$ and $a_{1},\ldots,a_{n}$ are in general position, the matrix $A_{I}$ is invertible. Now, $x\in\operatorname{Row}(A)$ is a linear combination of the rows of $A$ and $x_{i}=0$ for all $i\in I$. On the other hand, the rows of $A_{I}$ are linearly independent. It follows that $x$ is a trivial linear combination of the rows of $A$ and hence $x=0$. This shows that the columns $b_{j},j\in J$, are linearly independent. ∎
Under the additional assumption of general linear position, the face-dependence correspondence of Lemma A.2 simplifies and can be phrased in terms of positive spanning the whole space on the Gale-dual side. We are now ready to prove Lemma 2.2, which we restate for convenience.
Lemma A.4 (Faces vs. positive spanning subsets under general linear position).
Let $d^{\prime},d^{\prime\prime}\in\mathbb{N}$ and put $n:=d^{\prime}+d^{\prime\prime}$. Let $a_{1},\dots,a_{n}\in\mathbb{R}^{d^{\prime}}$ and $b_{1},\dots,b_{n}\in\mathbb{R}^{d^{\prime\prime}}$ be in linear Gale duality, and assume that both configurations are in general linear position. Let $I\subseteq[n]$ satisfy $I\leq d^{\prime}-1$. Then the following are equivalent:
  1. (i)
    $\mathop{\mathrm{pos}}\nolimits(a_{i}:i\in I)$ is a face of $C_{A}:=\mathop{\mathrm{pos}}\nolimits(a_{1},\dots,a_{n})$.
  2. (ii)
    $\mathop{\mathrm{pos}}\nolimits(b_{j}:j\in I^{c})=\mathbb{R}^{d^{\prime\prime}}$. Here $\II^{c}:=[n]\backslash I$ is the complement of $I$.
Proof.
We first consider the case $I=\varnothing$. Then $\mathop{\mathrm{pos}}\nolimits(a_{i}:i\in I)=\{0\}$, so (i) says that $\{0\}$ is a face of $C_{A}$. Recall that $a_{1},\ldots,a_{n}$ are in general linear position, in particular $a_{i}\neq 0$ for all $i\in[n]$. Hence $\{0\}$ is a face of $C_{A}$ if and only if there exists a face $F$ of $C_{A}$ containing none of the generators $a_{1},\dots,a_{n}$. Thus we may apply Lemma A.2 with $I=\varnothing$, which gives that (i) is equivalent to
$\mathop{\mathrm{pos}}\nolimits(b_{1},\ldots,b_{n})=\mathop{\mathrm{lin}}\nolimits(b_{1},\ldots,b_{n}).$
By Definition 2.1, we have $\mathop{\mathrm{lin}}\nolimits(b_{1},\ldots,b_{n})=\mathbb{R}^{d^{\prime\prime}}$. Thus (i) is equivalent to $\mathop{\mathrm{pos}}\nolimits(b_{1},\dots,b_{n})=\mathbb{R}^{d^{\prime\prime}}$, which is exactly (ii). Hence the lemma holds for $I=\varnothing$.
From now on assume $1\leq\#I\leq d^{\prime}-1$.
Claim. If $a_{1},\ldots,a_{n}$ are in general linear position and $I\subseteq[n]$ satisfies $1\leq\#I\leq d^{\prime}-1$, then
(A.4) $\{k\in[n]:a_{k}\in\mathop{\mathrm{pos}}\nolimits(a_{i}:i\in I)\}=I.$
Proof of the claim. Clearly $I$ is contained in the set on the left. Conversely, let $k\in[n]$ be such that $a_{k}\in\mathop{\mathrm{pos}}\nolimits(a_{i}:i\in I)$. Then $a_{k}\in\mathop{\mathrm{lin}}\nolimits(a_{i}:i\in I)$. If $k\notin I$, this gives a non-trivial linear dependence among the $I+1\leq d^{\prime}$ vectors $a_{i},i\in I\cup\{k\}$. Since $a_{1},\ldots,a_{n}$ are in general linear position, every set of at most $d^{\prime}$ vectors is linearly independent, a contradiction. This proves (A.4). ∎
[…]
$\mathop{\mathrm{pos}}\nolimits(b_{j}:j\in I^{c})=\mathop{\mathrm{lin}}\nolimits(b_{j}:j\in I^{c}).$
Since $I\leq d^{\prime}-1$, we have $(I^{c})=n-\#I\geq d^{\prime\prime}+1$. Since $b_{1},\ldots,b_{n}$ are in general linear position, $\mathop{\mathrm{lin}}\nolimits(b_{j}:j\in I^{c})=\mathbb{R}^{d^{\prime\prime}}$. Therefore $\mathop{\mathrm{pos}}\nolimits(b_{j}:j\in I^{c})=\mathbb{R}^{d^{\prime\prime}}$, which is (ii).
[…]
Acknowledgement
This paper benefited from many helpful and illuminating discussions with ChatGPT, which sharpened both the perspective and presentation; the work would have been impossible without these exchanges. Supported by the German Research Foundation under Germany’s Excellence Strategy EXC 2044/2 – 390685587, Mathematics Münster: Dynamics - Geometry - Structure and by the DFG priority program SPP 2265 Random Geometric Systems.

### v2

@@
-Random Polyhedral Cones: Distributional Results via Gale Duality
+Random Polyhedral Cones I: Distributional Results via Gale Duality
 Abstract.
 Let $U_{1},\ldots,U_{n}$ be independent random vectors uniformly distributed on the unit sphere $\mathbb{S}^{d-1}\subseteq\mathbb{R}^{d}$, where $n\geq d$, and consider the random polyhedral cone
-$\mathcal{W}_{n,d}:=\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{n})=\{\lambda_{1}U_{1}+\ldots+\lambda_{n}U_{n}:\lambda_{1}\geq 0,\ldots,\lambda_{n}\geq 0\}.$
+$\mathcal{W}_{n,d}:=\mathop{\mathrm{pos}}(U_{1},\ldots,U_{n})=\{\lambda_{1}U_{1}+\ldots+\lambda_{n}U_{n}:\lambda_{1}\geq 0,\ldots,\lambda_{n}\geq 0\}.$
 We establish several distributional results for $\mathcal{W}_{n,d}$ and the associated spherical polytope $\mathcal{W}_{n,d}\cap\mathbb{S}^{d-1}$. Our main contributions include:
@@
     Let $I_{1},I_{2}\subseteq[n]$ be such that $I_{1}\cup I_{2}=[n]$ and $\max(\#I_{1},\#I_{2})\leq d-1$.
-We prove that the “face events” $\{\,\mathop{\mathrm{pos}}\nolimits(U_{i}:i\in I_{r})\text{ is a face of }\mathcal{W}_{n,d}\,\}$, $r=1,2$, are independent.
+We prove that the “face events” $\{\,\mathop{\mathrm{pos}}(U_{i}:i\in I_{r})\text{ is a face of }\mathcal{W}_{n,d}\,\}$, $r=1,2$, are independent.
 A unifying ingredient is an explicit coupling producing i.i.d.
@@
 Key words and phrases:
-Stochastic geometry, geometric probability, random polyhedral cones, random spherical polytopes, solid angles, Gaussian projection, Gale duality, spherical Sylvester problem, $U$-statistics, Funk–Hecke formula, spherical harmonics
+Stochastic geometry, geometric probability, random polyhedral cones, random spherical polytopes, solid angles, Gale duality, spherical Sylvester problem, $U$-statistics, Funk–Hecke formula, spherical harmonics
 2020 Mathematics Subject Classification:
@@
 The random cones $\mathcal{W}_{n,d}$, $\mathcal{W}_{n,d}^{\circ}$, and related models have been studied by Cover and Efron [8], Donoho and Tanner [11], and Hug and Schneider [20]; see also [15, 21, 22, 25, 43] and Chapters 5–6 of the book by Schneider [44].
-One important result is Wendel’s formula [47] (see also [43, Theorem 8.2.1]), which states that
+One important result is Wendel’s formula [49] (see also [45, Theorem 8.2.1]), which states that
 (1.1) $p(n,d):=\mathbb{P}[\mathcal{W}_{n,d}=\mathbb{R}^{d}]=\frac{1}{2^{n-1}}\sum_{\ell=d}^{n-1}\binom{n-1}{\ell}$.
@@
 Gale duality (or the Gale transform) is a widely used tool in convex and discrete geometry.
-For detailed accounts of Gale duality and its applications, we refer to the books of Grünbaum [17, Section 5.4], Ziegler [50, Chapter 6],  Matoušek [33, Section 5.6], McMullen and Shephard [35, Chapter 3], Björner et al.
-[6, Section 3.4], and to the papers of McMullen [34], Shephard [46], and Grünbaum and Shephard [18].
+For detailed accounts of Gale duality and its applications, we refer to the books of Grünbaum [17, Section 5.4], Ziegler [52, Chapter 6],  Matoušek [33, Section 5.6], McMullen and Shephard [36, Chapter 3], Björner et al.
+[6, Section 3.4], Pineda Villavicencio [40, Section 2.14], McMullen [35, Sections 6B, 6C] and to the papers of McMullen [34], Shephard [48], and Grünbaum and Shephard [18].
 There are two versions of Gale duality: affine and linear.
@@
   2. (ii)
-    if $A\in\mathbb{R}^{d^{\prime}\times n}$ is the matrix with comumns $a_{1},\ldots,a_{n}$ and $B\in\mathbb{R}^{d^{\prime\prime}\times n}$ is the matrix with columns $b_{1},\ldots,b_{n}$, then every row of $A$ is orthogonal to every row of $B$.
+    if $A\in\mathbb{R}^{d^{\prime}\times n}$ is the matrix with columns $a_{1},\ldots,a_{n}$ and $B\in\mathbb{R}^{d^{\prime\prime}\times n}$ is the matrix with columns $b_{1},\ldots,b_{n}$, then every row of $A$ is orthogonal to every row of $B$.
 In matrix notation,
@@
 standard Gaussian points in $\mathbb{R}^{n}=\mathbb{R}^{d^{\prime}+d^{\prime\prime}}$.
-The linear Blaschke–Petkantschin formula [43, Theorem 7.2.1] states that for every nonnegative Borel function $h:(\mathbb{R}^{n})^{d^{\prime\prime}}\to[0,\infty)$,
+The linear Blaschke–Petkantschin formula [45, Theorem 7.2.1] states that for every nonnegative Borel function $h:(\mathbb{R}^{n})^{d^{\prime\prime}}\to[0,\infty)$,
 $\mathbb{E}h(X_{1},\ldots,X_{d^{\prime\prime}})=\int\displaylimits_{(\mathbb{R}^{n})^{d^{\prime\prime}}}\frac{{\rm e}^{-\frac{1}{2}(\|z_{1}\|^{2}+\ldots+\|z_{d^{\prime\prime}}\|^{2})}}{(2\pi)^{nd^{\prime\prime}/2}}\cdot h(z_{1},\ldots,z_{d^{\prime\prime}})\;\prod_{i=1}^{d^{\prime\prime}}\textup{d}z_{i}\\ =B\int\displaylimits_{G(n,d^{\prime\prime})}\int\displaylimits_{L}\frac{{\rm e}^{-\frac{1}{2}(\|z_{1}\|^{2}+\ldots+\|z_{d^{\prime\prime}}\|^{2})}}{(2\pi)^{nd^{\prime\prime}/2}}\cdot(\nabla_{d^{\prime\prime}}(z_{1},\ldots,z_{d^{\prime\prime}}))^{d^{\prime}}\cdot h(z_{1},\ldots,z_{d^{\prime\prime}})\;\prod_{i=1}^{d^{\prime\prime}}\lambda_{L}(\textup{d}z_{i})\;\nu_{n,d^{\prime\prime}}(\textup{d}L),$
-where $\nu_{n,d^{\prime\prime}}$ is the uniform probability distribution on $G(n,d^{\prime\prime})$, $\lambda_{L}$ is the Lebesgue measure on $L\in G(n,d^{\prime\prime})$, $\nabla_{d^{\prime\prime}}(z_{1},\ldots,z_{d^{\prime\prime}})$ denotes the $d^{\prime\prime}$-dimensional volume of the parallelepiped spanned by the vectors $z_{1},\ldots,z_{d^{\prime\prime}}$, and $B$ is a normalizing constant whose value can be found in [43, Eqn.
+where $\nu_{n,d^{\prime\prime}}$ is the uniform probability distribution on $G(n,d^{\prime\prime})$, $\lambda_{L}$ is the Lebesgue measure on $L\in G(n,d^{\prime\prime})$, $\nabla_{d^{\prime\prime}}(z_{1},\ldots,z_{d^{\prime\prime}})$ denotes the $d^{\prime\prime}$-dimensional volume of the parallelepiped spanned by the vectors $z_{1},\ldots,z_{d^{\prime\prime}}$, and $B$ is a normalizing constant whose value can be found in [45, Eqn.
 (7.8), p.
@@
 3.2. Expected face counts
-As a simple consequence of Theorem 2.8, we recover the following result of Donoho and Tanner [11, Theorem 1.6]. Recall that Wendel’s formula [47] (see also [43, Theorem 8.2.1]) states that
+As a simple consequence of Theorem 2.8, we recover the following result of Donoho and Tanner [11, Theorem 1.6]. Recall that Wendel’s formula [49] (see also [45, Theorem 8.2.1]) states that
 (3.2) $p(n,d)=\mathbb{P}[\mathcal{W}_{n,d}=\mathbb{R}^{d}]=\frac{1}{2^{n-1}}\sum_{\ell=d}^{n-1}\binom{n-1}{\ell}$.
@@
 Taking up a suggestion of Gale, Schneider [43] proposed to study random polytopes $P$ obtained by choosing their (affine) Gale transform at random.
-(Throughout this remark, we use the term “polytope” to refer to its combinatorial type; only the combinatorial type is canonically determined by the Gale transform.) The model of Schneider [41, p.
+(Throughout this remark, we use the term “polytope” to refer to its combinatorial type; only the combinatorial type is canonically determined by the Gale transform.) The model of Schneider [43, p.
 643] depends on an even probability measure $\phi$; in the present remark we take $\phi$ to be the uniform distribution on the unit sphere in $\mathbb{R}^{k}$.
@@
 Consequently, the random polytopes studied in Schneider [43] have the same distribution as the bounded cross-sections of Cover–Efron cones considered in [20].
-(As noted in [41, Remark on p.
+(As noted in [43, Remark on p.
 648], these two objects in particular have the same expected face numbers.)
@@
 Remark 4.5.
-In this section, we considered the cone $\mathcal{W}_{n,d}$ with $n=d$ only. Our methods do not extend to $n>d$. For example, it remains open to compute the variance of $\alpha(\mathcal{W}_{d+1,d})$.
+In this section, we considered the cone $\mathcal{W}_{n,d}$ with $n=d$ only.
+Our methods do not allow to compute the variance of $\alpha(\mathcal{W}_{n,d})$ for $n>d$.
+What we can show is that for $n\geq d$,
+(4.4) $\mathbb{E}[\alpha^{2}(\mathcal{W}_{n,d})]=\frac{\mathbb{E}[f_{1}^{2}(\mathcal{W}_{n+2,n+2-d})]-\mathbb{E}[f_{1}(\mathcal{W}_{n+2,n+2-d})]}{(n+2)(n+1)}.$
+Proof of (4.4).
+Let $k\in\mathbb{N}$.
+By Theorem 2.8, we may realize i.i.d.
+random vectors $U_{1},\ldots,U_{n+k}\sim\mathrm{Unif}(\mathbb{S}^{d-1})$ together the i.i.d.
+random vectors $V_{1},\ldots,V_{n+k}\sim\mathrm{Unif}(\mathbb{S}^{n+k-d-1})$ on a common probability space such that the duality (2.6) holds.
+We have
+$\mathbb{E}[\alpha^{k}(\mathcal{W}_{n,d})]$ $=\mathbb{P}\left[U_{n+1},\ldots,U_{n+k}\in-\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{n})\right]$
+$=\mathbb{P}\left[\mathop{\mathrm{pos}}\nolimits(U_{1},\ldots,U_{n},U_{n+i})=\mathbb{R}^{d}\ \forall i=1,\ldots,k\right].$
+Applying the duality (2.6), with $U_{i}$’s and $V_{j}$’s interchanged, to the right-hand side gives
+$\mathbb{E}[\alpha^{k}(\mathcal{W}_{n,d})]=\mathbb{P}\left[\mathop{\mathrm{pos}}\nolimits\bigl(\{V_{n+1},\ldots,V_{n+k}\}\setminus\{V_{n+i}\}\bigr)\text{ is a face of }\operatorname{pos}(V_{1},\ldots,V_{n+k})\ \forall i=1,\ldots,k\right].$
+If $k=2$, the condition on the right-hand side states that $\operatorname{pos}(V_{n+1})$ and $\operatorname{pos}(V_{n+2})$ are faces of $\operatorname{pos}(V_{1},\ldots,V_{n+2})$.
+So, using exchangeability,
+$\mathbb{E}[\alpha^{2}(\mathcal{W}_{n,d})]$ $=\mathbb{P}\left[\mathop{\mathrm{pos}}\nolimits(V_{n+1}),\mathop{\mathrm{pos}}\nolimits(V_{n+2})\in\mathcal{F}_{1}\left(\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{n+2})\right)\right]$
+$=\frac{1}{(n+2)(n+1)}\,\mathbb{E}\!\sum_{\begin{subarray}{c}i\neq j\\ i,j\in\{1,\ldots,n+2\}\end{subarray}}\mathbbm{1}_{\{\mathop{\mathrm{pos}}\nolimits(V_{i})\in\mathcal{F}_{1}(\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{n+2}))\}}\,\mathbbm{1}_{\{\mathop{\mathrm{pos}}\nolimits(V_{j})\in\mathcal{F}_{1}(\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{n+2}))\}}$
+$=\frac{1}{(n+2)(n+1)}\,\mathbb{E}\!\left[f_{1}^{2}\left(\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{n+2})\right)-f_{1}\left(\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{n+2})\right)\right],$
+and the proof is complete since $\mathop{\mathrm{pos}}\nolimits(V_{1},\ldots,V_{n+2})$ has the same law as $\mathcal{W}_{n+2,n+2-d}$. ∎
 4.3. Conjectures on high-dimensional random angles
@@
 The hyperplanes $U_{1}^{\perp},\ldots,U_{n}^{\perp}$ induce a conical tessellation of $\mathbb{R}^{d}$ into polyhedral cones.
-The number of cones is almost surely constant and equals $C(n,d)=2\sum_{\ell=0}^{d-1}\binom{n-1}{\ell}$, by the Steiner–Schläfli formula; see [43, Lemma 8.2.1].
+The number of cones is almost surely constant and equals $C(n,d)=2\sum_{\ell=0}^{d-1}\binom{n-1}{\ell}$, by the Steiner–Schläfli formula; see [45, Lemma 8.2.1].
 Choose one of these cones uniformly at random and denote it by $\mathcal{S}_{n,d}$.
@@
 The proof of Theorem 6.1 relies on the theory of $U$-statistics.
-In this section, we introduce the necessary notation and recall several classical results on $U$-statistics; see the books of Lee [29], Korolyuk and Borovskikh [27], and Serfling [45, Chapter 5] for further background.
+In this section, we introduce the necessary notation and recall several classical results on $U$-statistics; see the books of Lee [29], Korolyuk and Borovskikh [27], and Serfling [47, Chapter 5] for further background.
 Let $X_{1},X_{2},\ldots$ be i.i.d.
@@
 $\zeta_{c}:=\mathop{\mathrm{Var}}\nolimits g_{c}(X_{1},\ldots,X_{c})=\mathbb{E}\bigl[\widetilde{g}_{c}(X_{1},\ldots,X_{c})^{2}\bigr].$
-For $n\geq m$, the variance of $\mathbb{U}_{n}$ is given by the exact formula (see [29, Chapter 1, § 1.3, Theorem 3] or [45, Section 5.2.1, Lemma A])
+For $n\geq m$, the variance of $\mathbb{U}_{n}$ is given by the exact formula (see [29, Chapter 1, § 1.3, Theorem 3] or [47, Section 5.2.1, Lemma A])
 (6.7) $\mathop{\mathrm{Var}}\nolimits\mathbb{U}_{n}=\sum_{c=1}^{m}\frac{\binom{m}{c}\binom{n-m}{\,m-c\,}}{\binom{n}{m}}\;\zeta_{c}.$
@@
 constant.
-If $h$ is degenerate of order $c$, Equation (6.7) implies (see, e.g., [45, Section 5.3.4])
+If $h$ is degenerate of order $c$, Equation (6.7) implies (see, e.g., [47, Section 5.3.4])
 (6.8) $\mathop{\mathrm{Var}}\nolimits\mathbb{U}_{n}=\frac{\binom{m}{c}\binom{n-m}{m-c}}{\binom{n}{m}}\;\zeta_{c}\;+\;o(n^{-c})\sim\frac{\binom{m}{c}^{2}\,c!}{n^{c}}\;\zeta_{c},\qquad n\to\infty.$
-The next classical theorem characterizes the weak limit of the $U$-statisic $\mathbb{U}_{n}$ under degeneracy of order $2$; see [29, Section 3.2.2], [27, Chapter 4], [40, p.
-168] or [45, Section 5.5.2, p.
+The next classical theorem characterizes the weak limit of the $U$-statisic $\mathbb{U}_{n}$ under degeneracy of order $2$; see [29, Section 3.2.2], [27, Chapter 4], [42, p.
+168] or [47, Section 5.5.2, p.
 194].
@@
 To each closed convex cone $C\subseteq\mathbb{R}^{k}$ one can associate conic intrinsic volumes $\upsilon_{0}(C),\ldots,\upsilon_{k}(C)$, which are nonnegative and satisfy $\upsilon_{0}(C)+\cdots+\upsilon_{k}(C)=1$.
-We refer to [1, 2] and [42, Section 2.3] for their definition and properties.
+We refer to [1, 2] and [44, Section 2.3] for their definition and properties.
 Lemma 6.6 (Conditional expectation kernels).
@@
 In the conditional probability, the only random element is the random matrix $\mathbb{O}$.
-To compute the conditional probability, we invoke the conic kinematic formula whose proof can be found in [1, Corollary 5.2] or in [42, Theorem  4.3.5]; for applications see also [1].
+To compute the conditional probability, we invoke the conic kinematic formula whose proof can be found in [1, Corollary 5.2] or in [44, Theorem  4.3.5]; for applications see also [1].
 The formula states that for (deterministic) polyhedral cones $A,B\subseteq\mathbb{R}^{k}$ that are not both linear subspaces,
@@
 See also [24, Theorem 5.18] for an alternative approach.
-We sketch a proof of (6.15) for $n\in\{1,\dots,d\}$; for background on conic intrinsic volumes and hyperplane arrangements we refer to [42, Section 5.2].
+We sketch a proof of (6.15) for $n\in\{1,\dots,d\}$; for background on conic intrinsic volumes and hyperplane arrangements we refer to [44, Section 5.2].
 For each sign vector $\varepsilon=(\varepsilon_{1},\dots,\varepsilon_{n})\in\{\pm 1\}^{n}$, consider the random cone $C_{\varepsilon}:=\mathop{\mathrm{pos}}\nolimits(\varepsilon_{1}U_{1},\dots,\varepsilon_{n}U_{n})$.
@@
 On the other hand, with probability $1$, the cones $C_{\varepsilon}$ are precisely the chambers of a hyperplane arrangement in $\mathop{\mathrm{lin}}\nolimits(U_{1},\dots,U_{n})$ that has the same intersection poset as the coordinate hyperplane arrangement in $\mathbb{R}^{n}$, whose chambers are the $2^{n}$ orthants.
-Since the $k$th conic intrinsic volume of an orthant in $\mathbb{R}^{n}$ equals $\frac{1}{2^{n}}\binom{n}{k}$, the Klivans–Swartz formula [42, Theorem 5.1.4] yields $\sum_{\varepsilon\in\{\pm 1\}^{n}}\upsilon_{k}(C_{\varepsilon})=\binom{n}{k}$ a.s.
+Since the $k$th conic intrinsic volume of an orthant in $\mathbb{R}^{n}$ equals $\frac{1}{2^{n}}\binom{n}{k}$, the Klivans–Swartz formula [44, Theorem 5.1.4] yields $\sum_{\varepsilon\in\{\pm 1\}^{n}}\upsilon_{k}(C_{\varepsilon})=\binom{n}{k}$ a.s.
 Taking expectations gives $2^{n}\,\mathbb{E}\,\upsilon_{k}(\mathcal{W}_{n,d})=\sum_{\varepsilon\in\{\pm 1\}^{n}}\mathbb{E}\,\upsilon_{k}(C_{\varepsilon})=\binom{n}{k}$, which is precisely (6.15).
@@
 Proof.
-Let $U$ be uniformly distributed on $\mathbb{S}^{k-1}$. For all $v,w\in\mathbb{S}^{k-1}$ it is easy to check (see, e.g., [49, Lemma 6.7 on p. 144]) that
+Let $U$ be uniformly distributed on $\mathbb{S}^{k-1}$. For all $v,w\in\mathbb{S}^{k-1}$ it is easy to check (see, e.g., [51, Lemma 6.7 on p. 144]) that
 $\mathbb{E}\!\left[\operatorname{sign}\!\big(\langle v,U\rangle\big)\operatorname{sign}\!\big(\langle U,w\rangle\big)\right]=\frac{2}{\pi}\arcsin(\langle v,w\rangle).$
@@
 $\mathbb{P}\left[\left|\frac{f_{\ell}(\mathcal{W}_{n,d})}{\binom{n}{\ell}}-p(n-\ell,n-d)\right|\geq t\right]\leq 2{\rm e}^{-2t^{2}\lfloor n/(n-\ell)\rfloor}.$
-For the proof we need a result of Hoeffding [19] which can be found in the book of Serfling [45, Theorem A on p. 201] and in [39].
+For the proof we need a result of Hoeffding [19] which can be found in the book of Serfling [47, Theorem A on p. 201] and in [41].
 Lemma 6.19 (Hoeffding bound for $U$-statistics).
@@
 The next lemma makes precise how faces of the cone $\mathop{\mathrm{pos}}\nolimits(a_{1},\dots,a_{n})$ correspond to strictly positive dependent subsets (those for which $\mathop{\mathrm{pos}}\nolimits=\mathop{\mathrm{lin}}\nolimits$) in a Gale-dual configuration.
-It can be found in [46, Theorem 1 on p.
+It can be found in [48, Theorem 1 on p.
 167] or [34, 2A11 on p.